### v1

# Independence and mean sensitivity in minimal systems under group actions

Abstract.
Let $(X,G)$ be a topological dynamical system, where $G$ is an infinite countable discrete group. Suppose $(X,G)$ is minimal, has no essential $K$-IT-tuple, and one of the followings holds: (1) $(X,G)$ is incontractible; (2) $(X,G)$ is local Bronstein, and admits an invariant Borel probability measure. Then, $(X,G)$ is a regular $K^{\prime}$ to one extension of its maximal equicontinuous factor for some $11\leq K^{\prime}\leq K-$ 1. Especially, we give an affirmative answer to a conjecture proposed by Huang, Lian, Shao and Ye (J. Funct. Anal., 2021).
If $(X,G)$ is minimal where $G$ is amenable, under the local Bronstein condition, we establish parallel results regarding weak mean sensitivity and find that every weakly mean-sensitivity tuple is an IT-tuple.
Key words and phrases:
Independece, Minimality, Mean sensitivity
2020 Mathematics Subject Classification:
Primary: 37B05, 37A15
1. Introduction
1.1. Background
Throughout this paper, we always assume that $G$ is an infinite countable discrete group. By a topological dynamical system (tds for short) we mean a pair $(X,G)$, where $G$ acts on the compact metric space $X$ as a group of homeomorphisms.
In [5], Blanchard introduced the concept of entropy pairs as a tool to identify where entropy resides within a tds, and demonstrated that a tds has positive entropy if and only if it possesses essential entropy pairs. This approach, now widely known as local entropy theory (see [16, 24] for surveys), has become a valuable tool of entropy theory. Building on this foundation, Blanchard et al. [7] extended the concept by defining entropy pairs for invariant probability measures. Glasner and Weiss [23] further developed the notion of entropy tuples, while Huang and Ye [32] introduced entropy tuples for invariant probability measures. Concurrently, the study of sequence entropy tuples and tame systems evolved significantly [17, 22, 26, 27, 43, 46, 47], alongside the development of a theory of local complexity [6, 30, 31].
Kerr and Li extensively studied independence in [33, 34], uncovering its connection to (sequence) entropy tuples and tameness (i.e., the enveloping semigroup of this system is separable and Fréchet [20]). The concept of independence sets was introduced in [34, Definition 2.1] for a tds $(X,G)$ as follows. Given $K\in\mathbb{N}$, for a tuple $\mathcal{A}=(A_{1},\cdots,A_{K})$ of subsets of $X$, we say that a set $J\subset G$ is an independence set for $\mathcal{A}$ if for every nonempty finite subset $I\subset J$ and any function $\sigma:I\to\{1,2,\cdots,K\}$, we have
$\cap_{g\in I}g^{-1}A_{\sigma(g)}\neq\emptyset.$
A $K$-tuple $(x_{k})_{k=1}^{K}\in X^{K}$ is a $K$-IT-tuple (IT-pair for $K=2$) if for any open neighborhood $U_{k}$ of $x_{k}$ the tuple $(U_{1},U_{2},\cdots,U_{K})$ has an infinite independence set. If in addition, $x_{i}\neq x_{j}$ for any $i\neq j$, then $(x_{k})_{k=1}^{K}$ is called an essential $K$-IT-tuple.
Let $\pi_{eq}:X\to X_{eq}$ be the factor map to the maximal equicontinuous factor of $(X,G)$. Here the existence of $\pi_{eq}$ can refer to [10], and can be obtained in a constructive way via the regionally proximal equivalence relation [49]. For $K\in\mathbb{N}\cup\{\infty\}$, $\pi_{eq}$ is said to be almost $K$ to one if $\{y\in Y:|\pi^{-1}(y)|=K\}$ is a residual subset of $Y$. If $(X,G)$ is minimal, then $(X_{eq},G)$ is also minimal. Consequently, there exists a unique invariant measure $\nu_{eq}$. In this case, $\pi_{eq}$ is said to be regular $K$ to one if $\nu_{eq}(\{y\in X_{eq}:|\pi_{eq}^{-1}(y)|=K\})=1$.
It is intriguing that the number of elements in the preimage of $\pi_{eq}$ is closely related to the existence of IT-tuples. Glasner [17, Corollary 5.4] proved that if $(X,G)$ is a minimal and tame (i.e., has no essential IT-pair) tds with an invariant probability measure, then $\pi_{eq}$ is almost one to one. However, he left a question of whether this extension is regular one to one [17, Problem 5.7]. Subsequently, Fuhrmann, Glasner, Jäger, and Oertel [13, Theorem 1.2] gave an affirmative answer to this question. By establishing the existence of infinite independent sets on the Ellis semigroup of $(X_{eq},G)$, they found the infinite independent sets on $(X,G)$. In [29], Huang, Lian, Shao and Ye further developed the method in [13], and showed that if a minimal system is an almost finite to one extension of its maximal equicontinuous factor and has no infinite independent sets of length $k$ for some $k\geq 2$, then it has only finitely many ergodic measures. In the same paper, they proposed the following conjecture [29, Conjecture 2].
Conjecture 1.1.
Let $(X,G)$ be a minimal tds, where $G$ is amenable. If $\pi_{eq}$ is proximal and not almost one to one, then for each $K\geq 2$, there is an essential $K$-IT-tuple.
In this paper, we study the dynamics on hyperspace systems, and find a more direct method to search for infinite independent sets. As a corollary of our main result (Theorem 1.2), we provide an affirmative answer to Conjecture 1.1.
As another focus of this paper, we investigate sensitivity through localization theory. Motivated by local entropy theory, Ye and Zhang [52] introduced the concept of sensitive tuples. Li, Tu, and Ye [37] studied sensitivity in its mean form, while more recently, Li, Tu, Ye, and Yu [39, 41] explored the multi-version of mean sensitivity and its local representation, namely, mean sensitive tuples and mean sensitive tuples in the mean. Subsequently, Li, Liu, Tu, and Yu [40] demonstrated that for a minimal tds under integer group actions, every mean-sensitive tuple is an IT-tuple if $\pi_{eq}$ is almost one to one. They further conjectured that the almost one to one condition is unnecessary. As a corollary of our main results (see Theorem 1.5), we confirm the validity of their conjecture. In fact, we prove that this conjecture hold for a broader class of systems with actions by non-abelian groups. If, in addition, $\pi_{eq}$ is open, we show IT-tuples, IN-tuples, sequence entropy tuples and weakly mean-sensitive tuple all coincide.
1.2. Main results
In this subsection, we will state our main results and explain how they address the conjectures and questions mentioned above. In fact, we will also provide additional cases where our results can be applied effectively.
Let us begin with some definitions and notations. Let $(X,G)$ be a minimal tds.
  • (a)
    $(X,G)$ is said to be incontractible (see Glasner [18]) if for every $n\in\mathbb{N}$, $(X^{n},G)$ is Bronstein, i.e., the set of minimal points of $(X^{n},G)$ is dense in $X^{n}$.
  • (b)
    $(X,G)$ is said to satisfy the local Bronstein condition, or $(X,G)$ is local Bronstein, (see Auslander [3]) if, whenever $(x,x^{\prime})$ is a minimal point in $RP_{2}(X,G)$, there exist a sequence $\{(x_{n},x^{\prime}_{n})\}_{n\in\mathbb{N}}$ of minimal points in $X\times X$, and a sequence $\{g_{n}\}_{n\in\mathbb{N}}$ in $G$, such that
    $\lim_{n\to\infty}(x_{n},x^{\prime}_{n})=(x,x^{\prime})\text{ and }\lim_{n\to\infty}d(g_{n}x_{n},g_{n}x^{\prime}_{n})=0,$
    where
    $RP_{2}$ $(X,G)=\{((x_{1},x_{2})\in X^{2}:\forall\epsilon>0,~{}\exists x_{k}^{\prime}\in X\text{ and }g\in G\text{ with }$
    $d(x_{k},x_{k}^{\prime})\leq\epsilon\ (1\leq k\leq 2)\text{ and }d(gx^{\prime}_{k_{1}},gx_{k_{2}}^{\prime})\leq\epsilon\ (1\leq k_{1}\leq k_{2}\leq 2)\}.$
By a tds $(X,G)$, a nonempty set $A\subset X$ is called an IT-set if every tuple formed by points in $A$ is an IT-tuple. We now proceed to state our first main result.
Theorem 1.2.
Let $(X,G)$ be a minimal tds, and let $\pi_{eq}:X\to X_{eq}$ be the factor map to its maximal equicontinuous factor. Assume one of the following conditions holds,
  • (1)
    $(X,G)$ is incontractible;
  • (2)
    $(X,G)$ is local Bronstein, and admits an invariant Borel probability measure.
Then, there exists a dense $G_{\delta}$ subset $Z\subset X_{eq}$ with $\nu_{eq}(Z)=1$ such that $\pi_{eq}^{-1}(y)$ is an IT-set for every $y\in Z$.
Remark 1.3.
We note that both conditions in Theorem 1.2 ensure that the tds $(X,G)$ satisfies the local Bronstein condition. Indeed, it is straightforward to observe that if $(X^{2},G)$ is Bronstein, then $(X,G)$ is local Bronstein.
However, these additional conditions are essential. That is, simply assuming that $(X,G)$ is local Bronstein, is insufficient to establish our theorem. For example, let $\mathbb{T}=\mathbb{R}/\mathbb{Z}$. Define $T$ and $S$ on $\mathbb{T}$ by $Tx=x+\alpha\mod 1$ and $Sx=x^{2}$ for all $x\in\mathbb{T}$, where $\alpha$ is an irrational number. Let $G$ be the group generated by $T$ and $S$. Since $(\mathbb{T},T)$ is minimal, so is $(\mathbb{T},G)$. The maximal equicontinuous factor of $(\mathbb{T},G)$ is trivial, and so the fiber of $\pi_{eq}$ is $\mathbb{T}$. By the properties of $T$ and $S$, we deduce that the minimal points of $(\mathbb{T}\times\mathbb{T},G)$ lie in $\Delta_{\mathbb{T}}:=\{(x,x):x\in\mathbb{T}\}$, implying that $(\mathbb{T},G)$ satisfies the local Bronstein condition. Additionally, it can be verified that $(\mathbb{T},G)$ is c-ordered (see [22, Definition 2.6]). According to [22, Theorem 2.13] by Glasner and Megrelishvili, it follows that $(\mathbb{T},G)$ is null, and hence, it does not have any essential IT-pairs.
Now, we summarize some situations that meet the conditions in Theorem 1.2.
  1. (a)
    $(X,G)$ is a minimal tds, with $G$ as a virtually nilpotent group (in particular, any abelian group is a virtually nilpotent group).
    A nilpotent group has a lower central series that terminates in the trivial subgroup after finitely many steps. A virtually nilpotent group contains a finite-index nilpotent subgroup. Glasner [21] proved that in this case $(X,G)$ is incontractible. Since every virtually nilpotent group is an amenable group, it always admits an invariant probability measure. Thus, by Remark 1.3, $(X,G)$ also satisfies condition (2) in Theorem 1.2. By applying Theorem 1.2 in this situation, we give an affirmative answer to [44, Question 4.8], which asks whether any minimal tds $(X,\mathbb{Z})$ with finite maximal sequence entropy has finite type.
  2. (b)
    $(X,G)$ is a point distal tds.
    Following Veech [48], a tds $(X,G)$ is said to be point distal, if $(X,G)$ contains a distal point $x\in X$ whose orbit is dense in $X$. Here a point $x\in X$ is called a distal point if $\min_{g\in G}d(gx,gx^{\prime})>0$ for any $x^{\prime}\in X$ with $x^{\prime}\neq x$. Denote by $D(X,G)$ the set of distal points of $(X,G)$. Now we prove that each point distal system $(X,G)$ is incontractible.
    Let $(X,G)$ be a point distal system. It follows from Auslander-Ellis theorem that any distal point is a minimal point, and hence a point distal tds is minimal. Since the nonempty set $D(X,G)$ is $G$-invariant and $(X,G)$ is minimal, one has $D(X,G)$ is dense in $X$. By the definition of distal points, it is easy to check that $D(X,G)^{n}\subset D(X^{n},G)$, which implies that $D(X^{n},G)$ is dense in $X^{n}$ for each $n\in\mathbb{N}$. Using the fact that each distal point is a minimal point again, we have the set of minimal points of $(X^{n},G)$ is dense in $X^{n}$ for each $n\in\mathbb{N}$, and hence $(X,G)$ is incontractible.
  3. (c)
    ${(X,G)}$ is a minimal and almost finite to one extension of its maximal equicontinuous factor.
    It follows from [8, Theorem 3.16] that, if a minimal tds is an almost finite to one extension of its maximal equicontinuous factor, then it is a point distal tds. Thus, this case is a special case of situation (b). Huang, Lian, Shao and Ye [29] under the assumption that $\pi_{eq}$ is almost $N$ to one, and there is some integer $K\geq 2$ such that $(X,G)$ has almost no essential $K$-IT-tuples, proved that $(X,G)$ has no more than $N(K-1)^{N}$ ergodic measures. Recently, the upper bound was improved to $N(K-1)$ by Wang and the first two authors of this paper [42]. By applying Theorem 1.2 in this situation, we improve the upper bound to $(K-1)$ (see Corollary 4.2 (iv)).
    The upper bound of $(K-1)$ appears to be optimal. As shown in [13, Subsection 5.3], there exists a minimal system that has essential IT-pairs but no essential $3$-IT-tuples and exhibits two distinct ergodic probability measures.
  4. (d)
    $(X,G)$ is a minimal, proximal extension of its maximal equicontinuous factor, and admits an invariant probability measure (Establishing Conjecture 1.1).
    Recall that $\pi_{eq}$ is called a proximal extension if $\min_{g\in G}d(gx_{1},gx_{2})=0$ whenever $x_{1},x_{2}\in X$ with $\pi_{eq}(x_{1})=\pi_{eq}(x_{2})$. If a minimal point of $(X\times X,G)$ belongs to $RP_{2}(X,G)$ then $\pi_{eq}(x_{1})=\pi_{eq}(x_{2})$, implying $\min_{g\in G}d(gx_{1},gx_{2})=0$. In particular, $(X,G)$ is local Bronstein.
    Now we show Theorem 1.2 implies that Conjecture 1.1 holds. In fact, since $\pi_{eq}$ is proximal and not almost one to one, it follows from [29, Corollary 2.17] that $\pi_{eq}$ is infinite to one. Thus, Theorem 1.2 shows that for $\nu_{eq}$-a.e. $y\in X_{eq}$, $\pi_{eq}^{-1}(y)$ contains arbitrarily large essential IT-tuples, and hence Conjecture 1.1 holds. Notably, by Corollary 4.2 (iii), we prove a stronger result that $\pi_{eq}^{-1}(y)$ contains arbitrarily large essential IT-tuples for every $y\in X_{eq}$.
Now, let us turn to the second topic of our paper, mean sensitivity. To better illustrate, we first introduce the notion of weakly mean-sensitive tuple as follows.
Definition 1.4.
Let $(X,G)$ be a tds, where $G$ is amenable. A $K$-tuple $(x_{k})_{k=1}^{K}\in X^{K}$ is a weakly mean-sensitive $K$-tuple (weakly mean-sensitive pair for $K=2$) if for any open neighborhood $U_{k}$ of $x_{k}$, $k=1,2,\cdots,K$, there exists $\delta>0$ such that for any nonempty open subset $U$ of $X$,
$\overline{BD}(\{g\in G:(gU)\cap U_{k}\neq\emptyset\text{ for }1\leq k\leq K\})>\delta,$
where $\overline{BD}(A)$ is the upper Banach density of $A$ (see Section 2 for definition).
A set is called a weakly mean-sensitive set if every tuple composed of points within it is a weakly mean-sensitive tuple. Corresponding to Theorem 1.2 for IT-sets, we present the following results.
Theorem 1.5.
Let $(X,G)$ be a minimal tds, where $G$ is amenable. If $(X,G)$ satisfies the local Bronstein condition, then the following statements hold:
  1. (i)
    There exists a dense $G_{\delta}$ subset $Z\subset X_{eq}$ with $\nu_{eq}(Z)=1$ such that $\pi_{eq}^{-1}(y)$ is a weakly mean-sensitive set for every $y\in Z$.
  2. (ii)
    Assume $K\in\mathbb{N}$ and $\pi_{eq}$ is regular $K$ to one. Then, $(X,G)$ has essential weakly mean-sensitive $K$-tuple but has no essential weakly mean-sensitive $(K+1)$-tuple.
  3. (iii)
    Every weakly mean-sensitive set is an IT-set.
Remark 1.6.
Notice that the converse of Theorem 1.5 (iii) does not necessarily hold; specifically, there exist a minimal tds whose essential IT-pairs are not essential weakly mean-sensitive pairs. For example, Fuhrmann and Kwietniak [14, Section 3] constructed a class of minimal systems that are regular one to one extensions of its maximal equicontinuous factor, and possess essential IT-pairs. By Theorem 1.5 (ii), this class of systems has no essential weakly mean-sensitive pairs. Therefore, for a minimal tds, an IT-pair may not always be a weakly mean-sensitive pair.
The mean sensitive tuples was introduced by Li, Ye and Yu [41]. Let $(X,G)$ be a tds, where $G$ is amenable, and $\mathcal{F}$ be a Følner sequence of $G$. A $K$-tuple $(x_{k})_{k=1}^{K}\in X^{K}$ is a mean-sensitive $K$-tuple along $\mathcal{F}$ if for any open neighborhood $U_{k}$ of $x_{k}$ ($k=1,2,\cdots,K$), there exists $\delta>0$ such that for any non-empty open subset $U$ of $X$, there exist $y_{1},\cdots,y_{K}\in U$ such that
$\limsup_{n\to\infty}\frac{1}{|F_{n}|}|\{g\in F_{n}:gy_{k}\in U_{k}\text{ for }1\leq k\leq K\}|>\delta,$
where $|A|$ is the number of elements in a finite set $A$. It is straightforward to see that every mean-sensitive tuple is also a weakly mean-sensitive tuple. Li, Liu, Tu, and Yu [40] demonstrated that for a minimal tds $(X,G)$ with $\pi_{eq}$ being almost one to one, every mean-sensitive $K$-tuple along $\mathcal{F}$ is a $K$-IT-tuple when $G=\mathbb{Z}$ and =$\mathcal{F}=\{[1,n]\cap\mathbb{Z}\}_{n\in\mathbb{N}}$. As a direct application of Theorem 1.5, we establish a more general result.
Theorem 1.7.
Let $(X,G)$ be a minimal tds, where $G$ is amenable, and $\mathcal{F}$ be a Følner sequence of $G$. If $(X,G)$ satisfies the local Bronstein condition, then each mean-sensitive $K$-tuple along $\mathcal{F}$ is a $K$-IT-tuple for $K\in\mathbb{N}$ with $K\geq 2$.
For a tds $(X,G)$ and $K\geq 2$, a $K$-tuple $(x_{k})_{k=1}^{K}\in X^{K}$ is called a $K$-IN-tuple if, for any open neighborhoods $U_{k}$ of $x_{k}$, the tuple $(U_{1},U_{2},\ldots,U_{K})$ has arbitrarily long independent sets. A tuple is a sequence entropy tuple if and only if it is an IN-tuple [34, Theorem 5.9].
A factor map $\pi:(X,G)\to(Y,G)$ between two minimal tds is open if it maps open subsets of $X$ to open subsets of $Y$. Every distal extension is open (see [19]), and any factor map between minimal systems can be lifted to an open extension via almost one to one extensions [9, Chapter VI, Section 3].
Now, under the assumption in Theorem 1.5 and with $\pi_{eq}$ being open, we derive the following result, which is a direct consequence of Corollary 4.2 (ii) and Corollary 5.3 (ii).
Theorem 1.8.
Under assumption in Theorem 1.5 and with $\pi_{eq}$ being open, for $K\geq 2$, the weakly mean-sensitive $K$-tuple, the $K$-IT-tuple, the $K$-IN-tuple, and the $K$-sequence entropy tuple coincide.
Structure of this paper: In Section 2, we review basic concepts and prove some lemmas. In Section 3, we find the relation between a minimal tds and its hyperspace system, identifying a class of closed sets that satisfy the interior saturation property. In Section 4, we prove Theorem 1.2, while Section 5 focuses on the proof of Theorem 1.5.
2. Preliminaries
Throughout this paper, we denote by $\mathbb{N}$ the set of natural numbers. Let $X$ be a compact metric space with a metric $d$. We denote the interior, closure and boundary of a set $A$ in $X$ by $\text{int}(A)$, $\overline{A}$ and $\partial A$, respectively. For $x\in X$, a nonempty subset $V\subset X$ and $\epsilon>0$, denote $B^{X}_{\epsilon}(x)=\{x^{\prime}\in X:d(x,x^{\prime})<\epsilon\}$ and $B^{X}_{\epsilon}(V)=\{x^{\prime}\in X:d(V,x^{\prime})<\epsilon\}$. We denote the diameter of a set $A$ by $\text{diam}(A)$.
2.1. Subset of group
Let $G$ be a countable group. A set $\mathcal{S}\subset G$ is said to be a syndetic set of $G$ if there exists a finite set $K\subset G$ such that $Kg\cap\mathcal{S}\neq\emptyset\text{ for all }g\in G.$ A set $\mathcal{S}\subset G$ is said to be a thick set of $G$ if for any finite set $K\subset G$, there exists $g\in G$ such that $Kg\subset\mathcal{S}$. A set $\mathcal{S}\subset G$ is said to be a thickly syndetic set of $G$ if for any finite set $K\subset G$, the set $\{g\in G:Kg\subset\mathcal{S}\}$ is a syndetic set. Denote $\mathcal{F}_{s}(G)$, $\mathcal{F}_{t}(G)$, and $\mathcal{F}_{ts}(G)$ as the sets of syndetic, thick, and thickly syndetic sets of $G$, respectively. The following lemma is from [50, Proposition 3.2] and [4, Theorem 2.4].
Lemma 2.1.
Let $G$ be a countable group. Then, the following statements hold.
  • (1)
    If $\mathcal{S}_{1},\mathcal{S}_{2}\in\mathcal{F}_{ts}(G)$, then $\mathcal{S}_{1}\cap\mathcal{S}_{2}\in\mathcal{F}_{ts}(G)$.
  • (2)
    If $\mathcal{S}_{1}\in\mathcal{F}_{s}(G)$ and $\mathcal{S}_{2}\in\mathcal{F}_{t}(G)$, then $\mathcal{S}_{1}\cap\mathcal{S}_{2}\neq\emptyset$.
  • (3)
    If $\mathcal{S}_{1}\in\mathcal{F}_{s}(G)$ and $\mathcal{S}_{2}\in\mathcal{F}_{ts}(G)$, then $\mathcal{S}_{1}\cap\mathcal{S}_{2}\in\mathcal{F}_{s}(G)$.
Recall that a countable group $G$ is amenable (see [12]) if and only if there exists a sequence of finite, nonempty subsets =$\mathcal{F}=\{F_{n}\}_{n=1}^{\infty}$ of $G$, called a Følner sequence, such that
$\lim_{n\to\infty}\frac{|gF_{n}\Delta F_{n}|}{|F_{n}|}=0\text{ for every $g\in G$.}$
Given a subset $F$ of $G$, the upper and lower Banach density of $F$ are defined, respectively, by
$\overline{BD}(F)=\max_{\mathcal{F}}\limsup_{n\to\infty}\frac{|F_{n}\cap F|}{|F_{n}|}$
and
$\underline{BD}(F)=\min_{\mathcal{F}}\liminf_{n\to\infty}\frac{|F_{n}\cap F|}{|F_{n}|},$
where =$\mathcal{F}=\{F_{n}\}_{n=1}^{\infty}$ takes over all Følner sequences of $G$. It is clear that $\underline{BD}(F)\leq\overline{BD}(F)$ for any $F\subset G$. When a set is denoted with braces, for example $\{A\}$, we simply write $\underline{BD}\{A\}=\underline{BD}(\{A\})$ and $\overline{BD}\{A\}=\overline{BD}(\{A\})$.
2.2. Topological dynamical system
A tds $(X,G)$ is called minimal if $X$ contains no proper nonempty closed invariant subsets. We denote the orbit of $x\in X$ by $\text{Orb}(x,G)=\{gx:g\in G\}.$ A point $x\in X$ is called minimal if the system $(\overline{\text{Orb}(x,G)},G)$ is minimal. Denote by $Min(X,G)$ the set of all minimal points of $(X,G)$, which is nonempty. The closure of $Min(X,G)$ in $X$ is called the minimal center of $(X,G)$. A subset $M$ of $X$ is a minimal set if $M=\overline{\text{Orb}(x,G)}$ for some minimal point $x$ of $(X,G)$. For a subset $U$ of $X$ and $x\in X$, we denote
$N(x,U)=\{g\in G:gx\in U\}.$
The following results are well-known (see e.g. [15]).
Lemma 2.2.
Let $(X,G)$ be a tds. Then, the following statements hold:
  • (1)
    If $x\in X$ is a minimal point, then for any open neighborhood $V$ of $x$, $N(x,V)\in\mathcal{F}_{s}(G)$;
  • (2)
    If $\mu$ is an invariant probability measure of $(X,G)$, then for any measurable set $V\subset X$ with $\mu(V)>0$, $\{g\in G:\mu(gV\cap V)>0\}\in\mathcal{F}_{s}(G)$;
  • (3)
    Let $x\in X$, and $U\subset X$ be an open neighborhood of $\overline{Min(X,G)}$. Then, $N(x,U)\in\mathcal{F}_{ts}(G)$.
  • (4)
    Let $x\in X$, and let $M\subset\overline{\text{Orb}(x,G)}$ be a minimal set. Then, for any open neighborhood $U$ of $M$, $N(x,U)\in\mathcal{F}_{t}(G)$.
A factor map $\pi:X\to Y$ between two tds $(X,G)$ and $(Y,G)$ is a continuous onto map which intertwines the actions. In this case, we say that $(Y,G)$ is a factor of $(X,G)$, and $(X,G)$ is an extension of $(Y,G)$. The following results are well-known (see e.g. [2, p. 21, Proposition 14]).
Lemma 2.3.
Let $\pi:(X,G)\to(Y,G)$ be a factor map. Then the following statements hold:
  • (1)
    $\pi(Min(X,G))=Min(Y,G)$;
  • (2)
    If $(X,G)$ is minimal, then $\pi$ is semi-open, i.e., for any open subset $U$ of $X$, $\pi(U)$ has nonempty interior.
For $K\in\mathbb{N}\cup\{\infty\}$, a factor map $\pi:(X,G)\to(Y,G)$ is almost $K$ to one if $\{y\in Y:|\pi^{-1}(y)|=K\}$ is a residual subset of $Y$ (a residual set is a subset that contains a countable intersection of dense open sets). In particular, $\pi$ is said to be almost finite to one if this property holds for some $K<\infty$. It is well known that if $(X,G)$ is minimal, then $\pi:(X,G)\to(Y,G)$ is almost $K$ to one for $K=\min_{y\in Y}|\pi^{-1}(y)|$.
The notion of $K$-regionally proximal relation, introduced in [28] is defined for $K\geq 2$ by
$RP_{K}$ $(X,G)=\{(x_{i})_{i=1}^{K}\in X^{K}:\text{ for any }\epsilon>0\text{ there exist }x_{k}^{\prime}\in X\text{ and }g\in G$
[…]
Lemma 2.4.
Let $(X,G)$ be a minimal tds, and $(x_{1},x_{2},\cdots,x_{K})\in RP_{K}(X,G)$. Then, for any $\epsilon>0$ and nonempty open set $U\subset X$, there exists $h\in G$ and $x^{\prime}_{k}\in B_{\epsilon}^{X}(x_{k})$ for $k=1,2,\cdots,K$, such that $hx_{1}^{\prime},\cdots,hx_{K}^{\prime}\in U.$
Proof.
Since $(X,G)$ is minimal, $\{gU:g\in G\}$ is an open cover of $X$. Let $\gamma$ be a Lebesgue number of $\{gU:g\in G\}$. Since $(x_{1},x_{2},\cdots,x_{K})\in RP_{K}(X,G)$, there exists $h^{\prime}\in G$ and $x^{\prime}_{k}\in B_{\epsilon}^{X}(x_{k})$ for $k=1,2,\cdots,K$, such that
$d(h^{\prime}x^{\prime}_{k},h^{\prime}x^{\prime}_{k^{\prime}})<\gamma,1\leq k\leq k^{\prime}\leq K.$
There is $g\in G$ such that
$h^{\prime}x^{\prime}_{k}\in gU\text{ for }1\leq k\leq K.$
This implies $g^{-1}h^{\prime}x^{\prime}_{k}\in U$ for $1\leq k\leq K$. Then $h=g^{-1}h^{\prime}$ is as required. ∎
A tds $(X,G)$ is equicontinuous if for any $\epsilon>0$, there is $\delta>0$ such that whenever $x,y\in X$ with $d(x,y)<\delta$, then $d(gx,gy)<\epsilon$ for all $g\in G$. There is a smallest invariant equivalence relation $S_{eq}$ such that the quotient system $(X/S_{eq},G)$ is equicontinuous [11]. The equivalence relation $S_{eq}$ is called the equicontinuous structure relation, and the factor $(X/S_{eq},G)$ is called the maximal equicontinuous factor of $(X,G)$. We remark that $S_{eq}$ is the smallest closed $G$-invariant equivalence relation containing the $2$-regionally proximal relation $RP_{2}(X,G)$. The $2$-regional proximal relation $RP_{2}(X,G)$ is $G$-invariant, closed, symmetric, and reflective. It is easy to see that $RP_{2}(X,G)$ is an equivalence relation if and only if $S_{eq}=RP_{2}(X,G)$. In many cases, $RP_{2}(X,G)$ is an equivalence relation. For example, $(X,G)$ admits an invariant probability measure [45], or $(X,G)$ is local Bronstein [1].
Theorem 2.5 ([1, Theorem 8]).
Let $(X,G)$ be a minimal tds, and meets the local Bronstein condition. For $K\geq 2$, if $x_{1},x_{2},\cdots,x_{K}\in X$ satisfy $(x_{1},x_{k})\in RP_{2}(X,G)$ for $2\leq k\leq K$, then $(x_{1},x_{2},\cdots,x_{K})\in RP_{K}(X,G)$.
Remark 2.6.
The original condition in [1, Theorem 8] is a group-theoretic condition on the Ellis group of the system, which is weaker than the condition stated in Theorem 2.5; see [38, Remark 7.2] for a proof of Theorem 2.5.
Remark 2.7.
If the local Bronstein condition in Theorem 2.5 can be replaced by the assumption that $(X,G)$ admits an invariant Borel probability measure, then the requirement that $(X,G)$ is local Bronstein in Theorem 1.2 (2) becomes redundant. Notably, if this replacement holds, it would confirm the conjecture in [29, Conjecture 1], which states that for a minimal tds $(X,G)$ with an invariant probability measure and no essential IT-$K$-tuples for some $K\geq 2$, the map $\pi_{eq}$ must be almost finite to one.
2.3. Hyperspace system
Let $X$ be a compact metric space with metric $d$, and $2^{X}$ be the set of nonempty closed subsets of $X$ endowed with the Hausdorff metric $d_{H}$, which can be defined by
$d_{H}(A,B)=\max\{\max_{a\in A}d(a,B),\max_{b\in B}d(b,A)\},$
where $d(x,A)=\inf_{y\in A}d(x,y)$ for $x\in X$ and $A\subset X$. It is well known that $(2^{X},d_{H})$ is a compact metric space, as $X$ is a compact metric space. We say a sequence $\{A_{i}\}_{i=1}^{\infty}$ in $2^{X}$ converges to $A\in 2^{X}$, denoted by $\lim_{i\to\infty}A_{i}=A$, if
$\lim_{i\to\infty}d_{H}(A_{i},A)=0.$
The following results are well known (see e.g. [35, 36]).
Lemma 2.8.
Let $X,Y$ be compact metric spaces and $\pi:X\to Y$ be a continuous surjection. Then we have the following statements:
  1. (1)
    $\pi^{-1}:Y\to 2^{X},y\mapsto\pi^{-1}(y)$ is upper semi-continuous, and hence the continuity points of $\pi^{-1}$ form a dense $G_{\delta}$ subset of $Y$;
  2. (2)
    $\pi^{-1}$ is continuous if and only if $\pi$ is open.
Each tds $(X,G)$ induces a tds on $2^{X}$, known as the hyperspace system. More precisely, for $g\in G$, the action on $2^{X}$ is given by $gA=\{gx:x\in A\}$ for each $A\in 2^{X}$.
3. Interior saturation property
The aim of this section is to establish the relation between a minimal tds satisfying the conditions in Theorem 1.2 and its hyperspace system (see Proposition 3.4). However, for our purposes, we only need a corollary of this result, namely Lemma 3.9, which is applied in Section 4 and Section 5 to prove our main results. We believe that Theorem 1.2 will hold independent interest beyond the scope of this work.
3.1. Notation
In this subsection, we introduce some notations, and prove some basic properties about them.
Let $(X,G)$ be a minimal tds, and let $\pi_{eq}:(X,G)\to(X_{eq},G)$ be the factor map to the maximal equicontinuous factor of $(X,G)$. Then, $(X_{eq},G)$ is both minimal and uniquely ergodic. The unique ergodic measure is denoted by $\nu_{eq}$, and the metric on $X_{eq}$ is denoted by $d_{eq}$. Notice that using an equivalent metric we always assume that $(X_{eq},G)$ is an isometry, that is, $d_{eq}(y,y^{\prime})=d_{eq}(gy,gy^{\prime})$ for $y,y^{\prime}\in X_{eq},g\in G$.
Let
$\mathcal{X}=\overline{\{\pi_{eq}^{-1}(y):y\in X_{eq}\}}\subset 2^{X}$
and
$\pi_{\mathcal{X}}:\mathcal{X}\to X_{eq},~{}E\mapsto\pi_{eq}(E).$
Here $\mathcal{X}$ is a compact subspace of $2^{X}$ and $\pi_{\mathcal{X}}$ is well defined since $|\pi_{eq}(E)|=1$ for every $E\in\mathcal{X}$. We have the following description for the elements in $\mathcal{X}$.
Lemma 3.1.
Let $(X,G)$ be a minimal tds. If $\pi_{eq}$ is almost $N$ to one for some $N\in\mathbb{N}\cup\{\infty\}$, then $|E|\geq N$ for all $E\in\mathcal{X}$.
Proof.
Suppose, for contradiction, that there exists $E_{*}\in\mathcal{X}$ with $|E_{*}|=M<N$. Our aim is to show that $\{y\in X_{eq}:|\pi_{eq}^{-1}(y)|\leq M\}$ is residual in $X_{eq}$, contradicting the assumption that $\pi_{eq}$ is almost $N$ to one. This completes the proof of Lemma 3.1.
For $\gamma>0$ and $A\subset X$, denote
$s(A,\gamma)=\min\{m\in\mathbb{N}:\text{there exist }x_{1},\cdots,x_{m}\in X\text{ such that }A\subset\bigcup_{i=1}^{m}B^{X}_{\gamma}(x_{i})\}.$
Fix $\epsilon>0$ and $\delta>0$. Since $(X_{eq},G)$ is minimal and isometry, there exists a finite subset $F$ of $G$ such that
$\{gy:g\in F\}\text{ is }\delta\text{-dense in }X_{eq}\text{ for all }y\in X_{eq}.$ (3.1)
There is $\xi>0$ such that for any $x,x^{\prime}\in X$, $d(x,x^{\prime})<\xi\text{ implies }d(gx,gx^{\prime})<\epsilon\text{ for all }g\in F.$
Since $|E_{*}|=M$, there exists $y_{*}\in X_{eq}$ such that $s(\pi_{eq}^{-1}(y_{*}),\xi)\leq M$. Then, $s(\pi_{eq}^{-1}(gy_{*}),\epsilon)\leq M$ for all $g\in F.$ Thus, by (3.1), $Y_{\epsilon}:=\{y\in X_{eq}:s(\pi^{-1}_{eq}(y),\epsilon)\leq M\}$ is $\delta$-dense in $X_{eq}$. Noting that this holds for all $\delta>0$, one has $Y_{\epsilon}$ is dense in $X_{eq}$. Since $\pi_{eq}^{-1}$ is upper semi-continuous, $Y_{\epsilon}$ is an open subset of $X_{eq}$. Therefore, by the arbitrariness of $\epsilon>0$, $Y_{*}=\cap_{\epsilon>0}Y_{\epsilon}$ is a dense $G_{\delta}$ subset of $X_{eq}$. It is clear that
$|\pi_{eq}^{-1}(y)|\leq M<N\text{ for every }y\in Y_{*}.$
This implies $\pi_{eq}$ is almost $M$ to one, which is a contradiction. Hence $|E|\geq N$ for all $E\in\mathcal{X}$. ∎
Now, we construct a subsystem $(\mathcal{X}_{eq}^{\text{meas}},G)$ of $(2^{X},G)$ as follows. Put
$\mathcal{X}_{eq}^{\text{meas}}=\{E\in\mathcal{X}:\nu_{eq}(\pi_{\mathcal{X}}(B^{\mathcal{X}}_{\epsilon}(E)))>0,\forall\epsilon>0\}.$
It is clear that $\mathcal{X}_{eq}^{\text{meas}}$ is a closed and $G$-invariant subset of $2^{X}$. The following lemma shows that $\mathcal{X}_{eq}^{\text{meas}}$ contains full measure fibers of $X_{eq}$. In particular, it is nonempty, and hence $(\mathcal{X}_{eq}^{\text{meas}},G)$ is a subsystem of $(2^{X},G)$.
Lemma 3.2.
Let $(X,G)$ be a minimal tds and denote
$Y=\{y\in X_{eq}:\pi^{-1}_{eq}(y)\in\mathcal{X}_{eq}^{\text{meas}}\}.$
Then, $Y$ is a dense $G_{\delta}$ subset of $X_{eq}$ with $\nu_{eq}(Y)=1$.
Proof.
For $\epsilon>0$, denote by
$Y_{\epsilon}=\{y\in X_{eq}:\pi_{eq}^{-1}(y)\subset B^{X}_{\epsilon}(E)\text{ for some }E\in\mathcal{X}_{eq}^{\text{meas}}\}.$
Using the upper semi-continuity of $\pi_{eq}$, $Y_{\epsilon}$ is open for $\epsilon>0$. We claim that $Y=\cap_{n\in\mathbb{N}}Y_{1/n}$. First we show that $\cap_{n\in\mathbb{N}}Y_{1/n}\subset Y$. Let $y_{*}\in\cap_{n\in\mathbb{N}}Y_{1/n}$. There exists $E_{n}\in\mathcal{X}_{eq}^{\text{meas}}$ such that $\pi_{eq}^{-1}(y_{*})\subset B_{1/n}^{X}(E_{n})$ for $n\in\mathbb{N}$. Let $y_{n}=\pi_{\mathcal{X}}(E_{n})$. Then $\lim_{n\to\infty}y_{n}=y_{*}$. Passing to a subsequence, assume that $\lim_{n\to\infty}E_{n}=E_{*}$. Then,
$\pi_{eq}^{-1}(y_{*})\subset E_{*}\in\mathcal{X}_{eq}^{\text{meas}}$
and
$\pi_{\mathcal{X}}(E_{*})=\lim_{n\to\infty}\pi_{\mathcal{X}}(E_{n})=y_{*},$
which implies that $\pi^{-1}_{eq}(y_{*})=E_{*}\in\mathcal{X}_{eq}^{\text{meas}}$, thereby proving $y_{*}\in Y$. Due to the arbitrariness of $y_{*}\in\cap_{n\in\mathbb{N}}Y_{1/n}$, one has $\cap_{n\in\mathbb{N}}Y_{1/n}\subset Y$. Hence $Y=\cap_{n\in\mathbb{N}}Y_{1/n}$, as $Y\subset\cap_{n\in\mathbb{N}}Y_{1/n}$ is trivial. Therefore, $Y$ is a $G_{\delta}$ subset of $X$. It is easy to see that any continuity point of $\pi_{eq}^{-1}$ belongs to $Y$. Using Lemma 2.8, we know that $Y$ is dense in $X_{eq}$. Thus, $Y$ is a dense $G_{\delta}$ subset of $X_{eq}$.
On the other hand, for $\epsilon>0$, by Lusin’s theorem, we may find a measurable subset $Z_{\epsilon}$ of $X_{eq}$ with $\nu_{eq}(Z_{\epsilon})>1-\epsilon$ such that $\pi_{eq}^{-1}|_{Z_{\epsilon}}$ is continuous, and hence the support of $\nu_{eq}|_{Z_{\epsilon}}$ belongs to $Y$. We finish the proof of Lemma 3.2, as $\nu_{eq}(\cup_{\epsilon>0}Z_{\epsilon}))=1$. ∎
We now enumerate additional properties of $\mathcal{X}_{eq}^{\text{meas}}$ to facilitate its use in our subsequent work.
Lemma 3.3.
Let $(X,G)$ be a minimal tds. Then the following statements hold:
  • (1)
    If $\pi_{eq}$ is open, then $\mathcal{X}_{eq}^{\text{meas}}=\{\pi_{eq}^{-1}(y):y\in X_{eq}\}$;
  • (2)
    If $\pi_{eq}$ is regular $K$ to one, then $|E|\leq K$ for every $E\in\mathcal{X}_{eq}^{\text{meas}}$;
  • (3)
    If $\pi_{eq}$ is regular one to one, then $|E|=1$ for every $E\in\mathcal{X}_{eq}^{\text{meas}}$.
Proof.
(1) is from Lemma 2.8 (2); (2) is from the definition; (3) is from (2) and Lemma 3.1. ∎
3.2. Interior saturation property
For convenience, we say that a minimal tds $(X,G)$ has property ()(*)( ∗ ) if it meets the conditions in Theorem 1.2. That is, one of the followings holds:
  • (1)
    $(X,G)$ is incontractible;
  • (2)
    $(X,G)$ is local Bronstein, and admits an invariant Borel probability measure.
We say that a set $E\in 2^{X}$ has interior saturation property if
$\pi_{eq}^{-1}(\text{int}(\pi_{eq}(E)))\subset E.$ (3.2)
Denote by $\mathcal{E}_{isp}(X,G)$ the set of elements in $2^{X}$ having interior saturation property. The following result plays a crucial role in the proof of our main results.
Proposition 3.4.
Let $(X,G)$ be a minimal tds with property ()(*)( ∗ ), and let $U\subset X$ be a nonempty open subset. Denote by $M_{U}$ the minimal center of $(\overline{\text{Orb}(\overline{U},G)},G)$. Then,
$M_{U}\subset\mathcal{E}_{isp}(X,G).$
To establish Proposition 3.4, we begin by proving a weaker version.
Proposition 3.5.
Let $(X,G)$ be a minimal tds with property ()(*)( ∗ ), and let $U\subset X$ be a nonempty open subset. Assume $M\subset\overline{\text{Orb}(\overline{U},G)}$ is a minimal set. Then,
$M\subset\{E\in 2^{X}:\exists E^{\prime}\in\mathcal{X}^{meas}_{eq},E^{\prime}\subset E\}.$
Proof.
Let
=$\mathcal{M}=\{E\in 2^{X}:\exists E^{\prime}\in\mathcal{X}_{eq}^{meas},E^{\prime}\subset E\}.$
By the compactness and $G$-invariance of $\mathcal{X}_{eq}^{meas}$, it is straightforward to see that $\mathcal{M}$ is a compact and $G$-invariant subset of $2^{X}$. Combining this with the minimality of $M$, to prove Proposition 3.5, it suffices to show the existence of some $E_{0}\in M$ such that $E_{0}\in\mathcal{M}$. Furthermore, we only need to prove for any $\epsilon>0$, there exists $E_{\epsilon}\in M$ such that
$E_{\epsilon}\in\{E\in 2^{X}:\exists E^{\prime}\in\mathcal{X}^{meas}_{eq},E^{\prime}\subset B_{2\epsilon}^{X}(E)\}.$ (3.3)
Indeed, once this is established, by letting $\epsilon\searrow 0$ and invoking the compactness of $\mathcal{X}_{eq}^{meas}$ and $M$, we can obtain the existence of some $E_{0}\in M$ such that $E_{0}\in\mathcal{M}$.
We now prove (3.3) for a fixed $\epsilon>0$. Firstly, we show that
$\mathcal{E}_{\epsilon}^{X}=\{g\in G:\exists E^{\prime}\in\mathcal{X}^{meas}_{eq},E^{\prime}\subset B_{\epsilon}^{X}(g\overline{U})\}\in\mathcal{F}_{s}(G).$ (3.4)
Fix $E_{*}\in\mathcal{X}_{eq}^{meas}$, and let $y_{*}=\pi_{\mathcal{X}}(E_{*})$. Then,
$E_{*}\subset\pi_{eq}^{-1}(y_{*}).$ (3.5)
Let $\{x_{i}\}_{i=1}^{n}\text{ be an }({\epsilon}/{2})\text{-dense subset of }\pi^{-1}_{eq}(y_{*}).$ Due to Remark 1.3, both cases of property ()(*)( ∗ ) satisfy the local Bronstein condition. Using Lemma 2.4 and Theorem 2.5, we can find $\{\widetilde{x}_{i}\}_{i=1}^{n}\subset X$ and $g_{*}\in G$ such that
$g_{*}\widetilde{x}_{i}\in U\text{ and }d(x_{i},\widetilde{x}_{i})<\frac{\epsilon}{2}\text{ for }i=1,\cdots,n.$
There exists an open neighborhood $V_{i}$ of $\widetilde{x}_{i}$ for each $1\leq i\leq n$ such that
$g_{*}V_{i}\subset U,$ (3.6)
and
$d(z_{i},x_{i})<\frac{\epsilon}{2}\text{ whenever }z_{i}\in V_{i}.$
The later implies
$\pi_{eq}^{-1}(y_{*})\subset B^{X}_{\epsilon}(\{z_{1},\cdots,z_{n}\})\text{ whenever }(z_{1},\cdots,z_{n})\in V_{1}\times\cdots\times V_{n}.$ (3.7)
When the first case in property ()(*)( ∗ ) holds, we can find a minimal point $(x_{1}^{\prime},\cdots,x_{n}^{\prime})$ of $(X^{n},G)$ belonging to $V_{1}\times\cdots\times V_{n}$. By Lemma 2.2 (1),
$\mathcal{S}_{1}=\{g\in G:g(x_{1}^{\prime},\cdots,x_{n}^{\prime})\in V_{1}\times V_{2}\times\cdots\times V_{n}\}\in\mathcal{F}_{s}(G).$
Hence, for each $g\in\mathcal{S}_{1}$,
[…]
Thus, $\mathcal{E}_{\epsilon}^{X}$ is a syndetic set as it contains a syndetic set $\mathcal{S}_{1}g_{*}^{-1}$.
When the second case in property ()(*)( ∗ ) holds, we can find an ergodic probability measure $\mu$ of $(X,G)$. Since $(X,G)$ is minimal, it follows that $\mu$ has full support. Then,
$(\mu\times\mu\times\cdots\times\mu)(V_{1}\times V_{2}\times\cdots\times V_{n})>0.$
Thus, by Lemma 2.2 (2),
$\mathcal{S}_{2}$ =$=\{g\in G:g(V_{1}\times V_{2}\times\cdots\times V_{n})\cap(V_{1}\times V_{2}\times\cdots\times V_{n})\neq\emptyset\}\in\mathcal{F}_{s}(G).$
Hence, for $g\in\mathcal{S}_{2}$,
[…]
Thus, $\mathcal{E}_{\epsilon}^{X}$ is a syndetic set as it contains a syndetic set $\mathcal{S}_{2}g_{*}^{-1}$.
Now, we show that there exists $E_{\epsilon}\in M$ and $E_{\epsilon}^{\prime}\in\mathcal{X}_{eq}^{meas}$ such that $E_{\epsilon}^{\prime}\subset B_{2\epsilon}^{X}(E_{\epsilon})$. Since $M\subset\overline{\text{Orb}(\overline{U},G)}$ is a minimal set, by Lemma 2.2 (4),
$\mathcal{E}^{\mathcal{X}}_{\epsilon}=\{g\in G:g\overline{U}\in B^{\mathcal{X}}_{\epsilon}(M)\}\in\mathcal{F}_{t}(G).$ (3.8)
Combining (3.4), (3.8) and Lemma 2.1 (2), we can find $g_{\epsilon}\in\mathcal{E}^{\mathcal{X}}_{\epsilon}\cap\mathcal{E}_{\epsilon}^{X}$. Due to $g_{\epsilon}\in\mathcal{E}_{\epsilon}^{X}$, there exists $E_{\epsilon}^{\prime}\in\mathcal{X}_{eq}^{meas}$ such that $E_{\epsilon}^{\prime}\subset B_{\epsilon}^{X}(g_{\epsilon}\overline{U})$. Due to $g_{\epsilon}\in\mathcal{E}^{\mathcal{X}}_{\epsilon}$, there exists $E_{\epsilon}\in M$ such that $g_{\epsilon}\overline{U}\in B_{\epsilon}^{\mathcal{X}}(E_{\epsilon})$. Thus, $E_{\epsilon}^{\prime}\subset B_{2\epsilon}^{X}(E_{\epsilon}).$ Combined with the discussion at the beginning of the proof, this completes the proof of Proposition 3.5. ∎
To prove Proposition 3.4, we also need the following lemmas.
Lemma 3.6.
Let $(X,G)$ be a minimal tds. Then, the following statements hold:
  • (1)
    If $C\in 2^{X_{eq}}$ and $C_{*}\in\overline{\text{Orb}(C,G)}$, then $\text{diam}(C_{*})=\text{diam}(C)$.
  • (2)
    Let $C_{1},\cdots,C_{M}\in 2^{X}$ and $\{g_{n}\}_{n=1}^{\infty}$ be a sequence of $G$. If $\lim_{n\to\infty}g_{n}C_{m}=C_{m}^{\prime}$ for each $m\in\{1,2,\cdots,M\}$ and $\lim_{n\to\infty}g_{n}(\cup_{m=1}^{M}C_{m})=C^{\prime}$, then
    $C^{\prime}=\cup_{m=1}^{M}C_{m}^{\prime}.$
Proof.
(1) is straightforward from the assumption that $(X_{eq},G)$ is an isometry.
For (2), we define a map $\tau:(2^{X})^{M}\to 2^{X}$ by $\tau(A_{1},\cdots,A_{M})=\cup_{m=1}^{M}A_{m}$. By the definition of Hausdorff metric, we have for any $(A_{1},\cdots,A_{M}),(A_{1}^{\prime},\cdots,A_{M}^{\prime})\in(2^{X})^{M}$,
$d_{H}(\cup_{m=1}^{M}A_{m},\cup_{m=1}^{M}A_{m}^{\prime})\leq\max_{1\leq m\leq M}d_{H}(A_{m},A_{m}^{\prime}).$
which implies $\tau$ is continuous. Thus,
$C^{\prime}=\lim_{n\to\infty}g_{n}(\cup_{m=1}^{M}C_{m})$ $=\lim_{n\to\infty}g_{n}\tau(C_{1},\cdots,C_{M})$
$=\tau(\lim_{n\to\infty}g_{n}(C_{1}),\cdots,\lim_{n\to\infty}g_{n}(C_{M}))$
$=\tau(C_{1}^{\prime},\cdots,C_{m}^{\prime})=\cup_{m=1}^{N}C_{m}^{\prime}.$
The proof is complete. ∎
For any $F\in 2^{X}$, denote by
$V(F)=\{y\in X_{eq}:\exists E\in\mathcal{X}_{eq}^{meas},E\subset F,\pi_{\mathcal{X}}(E)=y\}.$ (3.9)
Lemma 3.7.
Let $(X,G)$ be a minimal tds. If $F\in 2^{X}$ satisfies
$V(F)=\pi_{eq}(F),$ (3.10)
then $F\in\mathcal{E}_{isp}(X,G)$.
Proof.
We fix $F\in 2^{X}$ satisfying (3.10), and prove that $F\in\mathcal{E}_{isp}(X,G)$. Denote by $Z_{0}$ the set of continuity points of the map $\pi_{eq}^{-1}$. Since $\pi_{eq}^{-1}$ is upper semi-continuous, by Lemma 2.8, $Z_{0}$ is a residual subset of $X_{eq}$. It is not difficult to check that $\pi_{\mathcal{X}}^{-1}(y)=\{\pi_{eq}^{-1}(y)\}$ for $y\in Z_{0}$. Since $F$ satisfies (3.10),
$\pi_{eq}^{-1}(Z_{0}\cap\text{int}(\pi_{eq}(F)))\subset F.$ (3.11)
We claim that
$\pi_{eq}^{-1}(Z_{0}\cap\text{int}(\pi_{eq}(F)))\text{ is dense in}\pi_{eq}^{-1}(\text{int}(\pi_{eq}(F))).$ (3.12)
If this is not true, we can find a nonempty open set $W\subset\pi_{eq}^{-1}(\text{int}(\pi_{eq}(F)))$ such that
$W\cap\pi_{eq}^{-1}(Z_{0}\cap\text{int}(\pi_{eq}(F)))=\emptyset.$ (3.13)
Thus,
$\pi_{eq}(W)\cap Z_{0}\cap\text{int}(\pi_{eq}(F))=\emptyset.$
Since $\pi_{eq}(W)\subset\text{int}(\pi_{eq}(F))$, it follows that $\pi_{eq}(W)\cap Z_{0}=\emptyset.$ This is impossible, as $\pi_{eq}$ being semi-open implies $\pi_{eq}(W)$ has nonempty interior, and $Z_{0}$ is dense in $X_{eq}$. Therefore, (3.12) holds.
Combining (3.11), (3.12) and the fact $F$ is a closed set, one obtains that
$\pi_{eq}^{-1}(\text{int}(\pi_{eq}(F)))\subset F.$
Thus, $F\in\mathcal{E}_{isp}(X,G)$. ∎
Lemma 3.8.
Let $(X,G)$ be a minimal tds. Then, the set
$\mathcal{V}=\{F\in 2^{X}:V(F)=\pi_{eq}(F)\}$
is a compact, $G$-invariant subset of $2^{X}$.
Proof.
The $G$-invariance of $\mathcal{V}$ is clear, and hence we only need to prove the compactness.
Let $E\in\mathcal{V}$, and suppose that there exists a sequence $\{E_{i}\}_{i=1}^{\infty}\subset 2^{X}$ with $V(E_{i})=\pi_{eq}(E_{i})$ for $i\in\mathbb{N}$, such that
$\lim_{i\to\infty}E_{i}=E.$ (3.14)
Thus,
$\lim_{i\to\infty}\pi_{eq}(E_{i})=\pi_{eq}(E).$
For any $y\in\pi_{eq}(E)$, there exists $y_{i}\in\pi_{eq}(E_{i})$ for $i\in\mathbb{N}$ such that
$\lim_{i\to\infty}y_{i}=y.$ (3.15)
By (3.9), for each $i\in\mathbb{N}$, there exists $E_{i}^{\prime}\in\mathcal{X}_{eq}^{meas}$ such that
$\pi_{\mathcal{X}}(E_{i}^{\prime})=y_{i},$ (3.16)
and
$E_{i}^{\prime}\subset E_{i}.$ (3.17)
Passing by a subsequence, we can assume that
$\lim_{i\to\infty}E_{i}^{\prime}=E^{\prime}\in\mathcal{X}_{eq}^{meas}.$ (3.18)
Combining (3.15), (3.16), and (3.18) yields $y=\pi_{\mathcal{X}}(E^{\prime})$, and from (3.14), (3.17), and (3.18), we deduce $E^{\prime}\subset E$. Thus $y\in V(E)$ and so $\pi_{eq}(E)\subset V(E)$. Since $V(E)\subset\pi_{eq}(E)$, we have $E\in\mathcal{V}$, thereby proving that $\mathcal{V}$ is a compact subset of $2^{X}$. ∎
Now we are able to prove Proposition 3.4.
Proof of Proposition 3.4.
By Lemma 3.7 and Lemma 3.8, to prove Proposition 3.4, it is sufficient to show that any $E_{*}\in Min(\overline{\text{Orb}(\overline{U},G)},G)$ satisfies $V(E_{*})=\pi_{eq}(E_{*})$. To do this, we only need to show that for any $\epsilon>0$, $V(E_{*})$ is $\epsilon$-dense in $\pi_{eq}(E_{*})$. Indeed, this implies $V(E_{*})=\pi_{eq}(E_{*})$ by taking $\epsilon\searrow 0$, since both $V(E_{*})$ and $\pi_{eq}(E_{*})$ are closed subset of $X_{eq}$.
Now, we begin proving the proof. Given $\epsilon>0$ and $E_{*}\in Min(\overline{\text{Orb}(\overline{U},G)},G)$, let $\{y_{1},\cdots,y_{I}\}\subset\pi_{eq}(U)$ be a finite $\epsilon/2$-dense subset of $\pi_{eq}(U)$. Then, $\pi_{eq}(\overline{U})\subset\cup_{i=1}^{I}B_{\epsilon/2}^{X_{eq}}(y_{i})$. For $i\in\{1,2,\cdots,I\}$, denote $B_{i}=\pi_{eq}^{-1}(B_{\epsilon/2}^{X_{eq}}(y_{i}))\cap U$, which is a nonempty open subset of $X$. Then,
$\text{diam}(\pi_{eq}(\overline{B_{i}}))<\epsilon,$ (3.19)
and
$\overline{U}=\cup_{i=1}^{I}\overline{B_{i}}.$ (3.20)
We consider a tds $(\overline{\text{Orb}((\overline{U},\overline{B_{1}},\cdots,\overline{B_{I}}),G)},G)$. Let
$p:(\overline{\text{Orb}((\overline{U},\overline{B_{1}},\cdots,\overline{B_{I}}),G)},G)\to(\overline{\text{Orb}(\overline{U},G)},G)$
be the factor map to the first coordinate, and let
$p_{i}:(\overline{\text{Orb}((\overline{U},\overline{B_{1}},\cdots,\overline{B_{I}}),G)},G)\to(\overline{\text{Orb}(\overline{B_{i}},G)},G)$
be the factor map to the $(i+1)$-th coordinate for each $i=1,\cdots,I$.
Since $E_{*}\in Min(\overline{\text{Orb}(\overline{U},G)},G)$, there exist a minimal set $M$ of $(\overline{\text{Orb}(\overline{U},G)},G)$ such that $E_{*}\in M$, and a minimal set $L$ of $(\overline{\text{Orb}((\overline{U},\overline{B_{1}},\cdots,\overline{B_{I}}),G)},G)$ such that $p(L)=M$.
Fix $(E_{*},E_{*,1},\cdots,E_{*,I})\in L.$ Then by Lemma 3.6 (2) and (3.20), we have
$E_{*}=\cup_{i=1}^{I}E_{*,i}.$ (3.21)
Since $E_{*,i}\in p_{i}(L)$ and $p_{i}(L)$ is a minimal set, it follows from Proposition 3.5 that there exists an element
$E_{i}^{\prime}\in\mathcal{X}_{eq}^{meas}\text{ such that }E_{i}^{\prime}\subset E_{*,i}\subset E_{*}\text{ for }i=1,\cdots,I.$ (3.22)
Consider the tds $(Y,G)=((\pi_{eq}\times\cdots\times\pi_{eq})(\overline{\text{Orb}((\overline{U},\overline{B_{1}},\cdots,\overline{B_{I}}),G)}),G)$. Due to the fact that $\pi_{eq}$ is a factor map, it follows that
$Y=\overline{\text{Orb}((\pi_{eq}(\overline{U}),\pi_{eq}(\overline{B_{1}}),\cdots,\pi_{eq}(\overline{B_{I}})),G)}.$
Thus, $(\pi_{eq}(E_{*}),\pi_{eq}(E_{*,1}),\cdots,\pi_{eq}(E_{*,I}))\in Y$. By (3.19) and Lemma 3.6 (1),
$\text{diam}(\pi_{eq}(E_{*,i}))=\text{diam}(\pi_{eq}(B_{i}))<\epsilon\text{, for $i=1,\cdots,I$.}$ (3.23)
By (3.21), $\pi_{eq}(E_{*})=\cup_{i=1}^{I}\pi_{eq}(E_{*,i})$. Combining this with (3.22) and (3.23),
$\{\pi_{\mathcal{X}}(E_{i}^{\prime}):i=1,\cdots,I\}\text{ is }\epsilon\text{-dense in }\pi_{eq}(E_{*}).$
By (3.22), each $\pi_{\mathcal{X}}(E_{i}^{\prime})\in V(E_{*})$ and so $V(E_{*})$ is $\epsilon$-dense in $\pi_{eq}(E_{*})$. Thus, Proposition 3.4 follows from the arguments at beginning. ∎
As a corollary of Proposition 3.4, we present the following lemma, which will be useful in Section 4 and Section 5.
Lemma 3.9.
Let $(X,G)$ be a minimal tds with property ()(*)( ∗ ), and let $U\subset X$ be a nonempty open set with $B^{X_{eq}}_{r}(y_{*})\subset\pi_{eq}(\overline{U})$ for some $r>0$ and $y_{*}\in X_{eq}$. Then, the following statements hold:
  • (1)
    For every pair $(y,E)\in\overline{\text{Orb}((y_{*},\overline{U}),G)}$, one has $B^{X_{eq}}_{r}(y)\subset\pi_{eq}(E)$.
  • (2)
    For every $\epsilon\in(0,r)$, one has
    $\{g\in G:\pi_{eq}^{-1}(B^{X_{eq}}_{r-\epsilon}(gy_{*}))\subset\overline{B^{X}_{\epsilon}(gU)}\}\in\mathcal{F}_{ts}(G).$
Proof.
For (1), let $(y,E)\in\overline{\text{Orb}((y_{*},\overline{U}),G)}$. There exists a sequence $\{g_{n}\}_{n=1}^{\infty}$ of $G$ such that $\lim_{n\to\infty}g_{n}y_{*}=y$ and $\lim_{n\to\infty}g_{n}\overline{U}=E$. Since $\pi_{eq}$ is continuous, it follows that $\lim_{n\to\infty}g_{n}\pi_{eq}(\overline{U})=\pi_{eq}(E)$. Note that $B^{X_{eq}}_{r}(y_{*})\subset\pi_{eq}(\overline{U})$ and $(X_{eq},G)$ is an isometry. One has $B^{X_{eq}}_{r}(g_{n}y_{*})\subset g_{n}\pi_{eq}(\overline{U})$ for each $n\in\mathbb{N}$, and thus, by letting $n\to\infty$, $B^{X_{eq}}_{r}(y)\subset\pi_{eq}(E)$.
For (2), given $\epsilon\in(0,r)$, let
$\mathcal{U}_{\epsilon}=\{(y,E)\in X\times 2^{X}:\min_{(y^{\prime},E^{\prime})\in M_{y_{*},U}}\{d_{eq}(y,y^{\prime})+d_{H}(E,E^{\prime})\}<\epsilon\},$
where $M_{y_{*},U}$ is the minimal center of $(\overline{\text{Orb}((y_{*},\overline{U}),G)},G)$. Then, $\mathcal{U}_{\epsilon}$ is an open neighborhood of $M_{y_{*},U}$. By Lemma 2.2 (3), one has
$\{g\in G:g(y_{*},\overline{U})\in\mathcal{U}_{\epsilon}\}\in\mathcal{F}_{ts}(G).$ (3.24)
Next, we will prove that for each $(y,E)\in\mathcal{U}_{\epsilon}$, the inclusion $\pi_{eq}^{-1}(B^{X_{eq}}_{r-\epsilon}(y))\subset\overline{B^{X}_{\epsilon}(E)}$ holds. This directly implies
$\{g\in G:\pi_{eq}^{-1}(B^{X_{eq}}_{r-\epsilon}(gy_{*}))\subset\overline{B^{X}_{\epsilon}(gU)}\}\supset\{g\in G:g((y_{*},\overline{U}))\in\mathcal{U}_{\epsilon}\}.$
By combining this with (3.24), the proof is complete. Indeed, fixing $(y,E)\in\mathcal{U}_{\epsilon}$, there exists $(y^{\prime},E^{\prime})\in M_{y_{*},U}$ such that $d_{eq}(y,y^{\prime})+d_{H}(E,E^{\prime})<\epsilon$. Thus, by Lemma 3.9 (1), one has
$\pi_{eq}^{-1}(B^{X_{eq}}_{r-\epsilon}(y))\subset\pi_{eq}^{-1}(B^{X_{eq}}_{r}(y^{\prime}))\subset\pi_{eq}^{-1}(\text{int}(\pi_{eq}(E^{\prime}))).$ (3.25)
Note that $E^{\prime}$ belongs to the minimal center of $(\overline{\text{Orb}(\overline{U},G)},G)$. Then by Proposition 3.4, We know that $E^{\prime}$ has interior saturation property, which together with (3.25), implies that
$\pi_{eq}^{-1}(B^{X_{eq}}_{r-\epsilon}(y))\subset E^{\prime}\subset B^{X}_{\epsilon}(E).$
The proof is completed. ∎
4. IT-set
In this section, we prove Theorem 1.2. To achieve this, we establish a stronger result, Theorem 4.1. The key lies in discovering a new approach (Claim 4.5) for constructing infinite independent sets.
4.1. Main result and corollary
Recall that
$\mathcal{X}_{eq}^{\text{meas}}=\{E\in\mathcal{X}:\nu_{eq}(\pi_{\mathcal{X}}(B^{\mathcal{X}}_{\epsilon}(E))>0,\forall\epsilon>0\},$
where $\mathcal{X}=\overline{\{\pi_{eq}^{-1}(y):y\in X_{eq}\}}\text{ and }\pi_{\mathcal{X}}:\mathcal{X}\to X_{eq},~{}E\mapsto\pi_{eq}(E)$. In Lemma 3.2, we have proven that the set $\{y\in X_{eq}:\pi_{eq}^{-1}(y)\in\mathcal{X}_{eq}^{meas}\}$ is a dense $G_{\delta}$ subset of $X_{eq}$ with full $\nu_{eq}$-measure. Thus, Theorem 1.2 is a direct corollary of the following theorem.
Theorem 4.1.
Let $(X,G)$ be a minimal tds with property ()(*)( ∗ ). Then, every $E_{*}\in\mathcal{X}_{eq}^{\text{meas}}$ is an IT-set.
The proof of Theorem 4.1 is intricate; therefore, we defer its presentation to Subsection 4.2. For now, we assume its validity to derive some corollaries.
Corollary 4.2.
Let $(X,G)$ be a minimal tds with property ()(*)( ∗ ). Then, the following statements hold:
  1. (i)
    If $(X,G)$ has no essential $K$-IT-tuple for some $K\geq 2$, then $(X,G)$ is regular $K^{\prime}$ to one for some $11\leq K^{\prime}\leq K-$ 1.
  2. (ii)
    If $\pi_{eq}$ is open, then $\pi_{eq}^{-1}(y)$ is an IT-set for every $y\in X_{eq}$.
  3. (iii)
    Let $N\in\mathbb{N}\cup\{\infty\}$ and $K\in\mathbb{N}$ with $2\leq K\leq N$. If $\pi_{eq}$ is almost $N$ to one, then for every $y\in X_{eq}$, $\pi_{eq}^{-1}(y)$ contains essential $K$-IT-tuples.
  4. (iv)
    If $(X,G)$ has no essential $K$-IT-tuples for some $K\geq 2$, then $|\mathcal{M}^{e}(X,G)|\leq K-1$.
Proof.
(i) follows directly from Theorem 4.1 and Lemma 3.2. Since $\pi_{eq}$ is open, by Lemma 3.3 (2), $\mathcal{X}_{eq}^{meas}=\{\pi_{eq}^{-1}(y):y\in X_{eq}\}$. Hence, (ii) follows from Theorem 4.1. (iii) is established using Theorem 4.1 and Lemma 3.1.
Now we prove (iv) by contradiction. Suppose that there exist distinct ergodic probability measures $\mu_{1},\mu_{2},$ $\cdots,\mu_{K}$ of $(X,G)$. Then there exist pairwise disjoint measurable subsets $V_{1},V_{2},\cdots,V_{K}$ of $X$ such that
$\mu_{k}(V_{k})\geq 1-\frac{1}{K+1}\text{ for }k=1,2,\cdots,K.$
Then,
$\nu_{eq}\left(\cap_{k=1}^{K}\pi_{eq}(V_{k})\right)\geq\frac{1}{K+1}>0.$
By Theorem 4.1 and Lemma 3.2, there exists $y_{*}\in\bigcap_{k=1}^{K}\pi_{eq}(V_{k})$ such that $\pi_{eq}^{-1}(y_{*})$ is an IT-set. Since $V_{1},V_{2},\cdots,V_{K}$ are pairwise disjoint, it follows that $|\pi_{eq}^{-1}(y_{*})|\geq K$. Take distinct points $x_{1},\cdots,x_{K}\in\pi_{eq}^{-1}(y_{*})$. Then, $(x_{1},\cdots,x_{K})$ is an essential $K$-IT-tuple since $\pi_{eq}^{-1}(y_{*})$ is an IT-set, which is a contradiction. We finish the proof. ∎
4.2. Proof of Theorem 4.1
Given a tds $(X,G)$, the Ellis semigroup $E(X)$ associated to $(X,G)$ is defined as the closure of $\{x\mapsto gx:g\in G\}\subset X^{X}$ in the product topology, where the semigroup operation is given by the composition. On $E(X)$, we may consider the tds given by $E(X)\ni s\mapsto gs$ for each element $g\in G$.
We now summarize some properties of the tds $(E(X),G)$, as detailed in [2, pp. 52-53], [51, Theorems 5.4 and 6.20] and [25, Theorem 15.13] for items (a)-(c). Property (d) follows directly from the fact that $E(H)$ is a group.
Theorem 4.3.
Suppose $H$ is a compact metric space and $(H,G)$ is minimal and equicontinuous. Then we have the following statements:
  1. (a)
    $E(H)$ is a compact metrisable topological group, and the left Haar measure of $E(H)$ coincides with the right Haar measure of $E(H)$.
  2. (b)
    $G$ as a group of homeomorphisms on $H$, can be seen as a subgroup of $E(H)$, and $E(H)$ can be seen as the closure of $G$. Hence, $(E(H),G)$ is both minimal and uniquely ergodic.
  3. (c)
    $(H,G)$ is a factor of $(E(H),G)$, with an open factor map $\pi_{H}$ is given by
    $\pi_{H}:E(H)\to H,t\mapsto th,\text{ for any fixed $h\in H$}.$
  4. (d)
    For $Z,W\subset E(H)$, $gW\cap Z\neq\emptyset$ if and only if $g\in ZW^{-1}=\{zw^{-1}:z\in Z,w\in W\}$.
Lemma 4.4.
Let $H$ be a compact metrisable topological group, and let $\nu_{H}$ be the Haar measure on $H$. Assume $A$ is a measurable subset of $H$ and $h_{1},h_{2},\cdots,h_{M}\in H$ and $\kappa\geq 0$. If $\nu_{H}(\cap_{m=1}^{M}Ah_{m}^{-1})>\kappa$ then there exist an open neighborhood $V_{m}$ of $h_{m}$ for all $m=1,2,\cdots,M$ such that
$\nu_{H}(\cap_{m=1}^{M}Ah_{m}^{\prime-1})>\kappa\text{ whenever }h_{m}^{\prime}\in V_{m}.$
Proof.
By assumption, there exists a compact set $A_{*}\subset\cap_{m=1}^{M}Ah_{m}^{-1}$ with $\nu_{H}(A_{*})=\kappa_{*}>\kappa$. Thus, there exists an open neighborhood $W$ of $A_{*}$ such that $\nu_{H}(W\setminus A_{*})<\frac{1}{M}(\kappa_{*}-\kappa)$. Choose an open neighborhood $U$ of $e_{H}$ such that $A_{*}U\subset W$, where $e_{H}$ is the identity of $H$. Then,
$\nu_{H}((A_{*}h)\setminus A_{*})<\frac{1}{M}(\kappa_{*}-\kappa),\text{ for all }h\in U.$ (4.1)
Since $V_{m}=Uh_{m}$ is an open neighborhood of $h_{m}$, and for any $h_{m}^{\prime}\in V_{m}$, $m\in\{1,2,\cdots,M\}$, one has that
$\nu_{H}(\cap_{m=1}^{M}Ah_{m}^{\prime-1})$ $=\nu_{H}(\cap_{m=1}^{M}(Ah_{m}^{-1}(h_{m}^{\prime}h_{m}^{-1})^{-1}))$
$\geq\nu_{H}(\cap_{m=1}^{M}(A_{*}(h_{m}^{\prime}h_{m}^{-1})^{-1})$
$\geq\nu_{H}(A_{*})-\sum_{1\leq m\leq M}\nu_{H}(A_{*}\setminus A_{*}(h_{m}^{\prime}h_{m}^{-1})^{-1})$
$=\nu_{H}(A_{*})-\sum_{1\leq m\leq M}\nu_{H}((A_{*}h_{m}^{\prime}h_{m}^{-1})\setminus A_{*})$
$>\kappa.$
Here the last inequality is from the fact that $h_{m}^{\prime}h_{m}^{-1}\in U$ and (4.1). ∎
Now we are able to prove Theorem 4.1.
Proof of Theorem 4.1.
Given any $E_{*}\in\mathcal{X}_{eq}^{\text{meas}}$, let $K\in\mathbb{N}$ and $x_{1},x_{2},\cdots,x_{K}\in E_{*}$. Now we show for any $\epsilon_{*}>0$, there exists an infinite independence set for $(B^{X}_{\epsilon_{*}}(x_{1}),\cdots,B^{X}_{\epsilon_{*}}(x_{K}))$. This finishes the proof of Theorem 4.1, by the arbitrariness of $x_{1},x_{2},\cdots,x_{K}\in E_{*}$.
Given $\epsilon_{*}>0$, let $Z_{*}=\pi_{\mathcal{X}}(B_{\epsilon_{*}/2}^{\mathcal{X}}(E_{*}))$. Since $E_{*}\in\mathcal{X}_{eq}^{\text{meas}}$, it follows that $\nu_{eq}(Z_{*})>0$. Moreover, as $E_{*}\subset B_{\epsilon_{*}/2}^{X}(E)\text{ for }E\in B_{\epsilon_{*}/2}^{\mathcal{X}}(E_{*})$ and $E\subset\pi_{eq}^{-1}(\pi_{\mathcal{X}}(E))$ for every $E\in\mathcal{X}$, we have
$E_{*}\subset B_{\epsilon_{*}/2}^{X}(\pi^{-1}_{eq}(y))\text{ for }y\in Z_{*}.$ (4.2)
Let $E(X_{eq})$ be the Ellis semigroup of $(X_{eq},G)$. By Theorem 4.3, $E(X_{eq})$ is a compact metrisable topological group and there is an open factor map
$\tilde{\pi}:(E(X_{eq}),\tilde{\nu},G)\to(X_{eq},\nu_{eq},G),$
where $\tilde{\nu}$ is the Haar measure on $E(X_{eq})$, and $\nu_{eq}=\tilde{\nu}\circ\tilde{\pi}^{-1}$. The relation is provided in the following diagram:
${(X,G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{eq}}$${(E(X_{eq}),\tilde{\nu},G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\pi}}$${(X_{eq},\nu_{eq},G).}$
Set $\tilde{Z}_{*}=\tilde{\pi}^{-1}(Z_{*})$, and we have the following claim.
Claim 4.5.
Given $M\in\mathbb{N}$, assume nonempty open subsets $W_{1},\cdots,W_{M}$ of $X$, and $\{h_{m}\}_{1\leq m\leq M}\subset E(X_{eq})$ satisfy the following conditions:
  • (P1)
    $\tilde{\pi}(h_{m})\in\pi_{eq}(W_{m})$ for $m=1,2,\cdots,M$;
  • (P2)
    $\tilde{\nu}(\cap_{m=1}^{M}(\tilde{Z}_{*}h_{m}^{-1}))>0.$
Then, for any finite set $F\subset G$, there exists $g_{*}\in G\setminus F$, and nonempty open sets
$W_{m,k}\subset W_{m}\text{ with }g_{*}W_{m,k}\subset B_{\epsilon_{*}}^{X}(x_{k})$
for $1\leq m\leq M,1\leq k\leq K$, and $\{h_{m,k}\}_{1\leq m\leq M,1\leq k\leq K}\subset E(X_{eq})$ such that
  • (P1’)
    $\tilde{\pi}(h_{m,k})\in\pi_{eq}(W_{m,k})$ for $m=1,2,\cdots,M$, $k=1,2,\cdots,K$;
  • (P2’)
    $\tilde{\nu}(\cap_{m=1}^{M}\cap_{k=1}^{K}(\tilde{Z}_{*}h_{m,k}^{-1}))>0.$
The proof of Claim 4.5 is complex, so we will present it later. For now, we assume that this claim holds to complete the proof of Theorem 4.1.
Put $W_{0}=X$ and take $h_{0}\in E(X_{eq})$. It is clear that $\tilde{\pi}(h_{0})\in\pi_{eq}(W_{0})$, and $\tilde{\nu}(\tilde{Z}_{*}h_{0}^{-1})=\tilde{\nu}(\tilde{Z}_{*})>0$. Thus, applying Claim 4.5, we can find $g_{1}\in G$, and nonempty open sets $W_{k}\subset W_{0}\text{ with }g_{1}W_{k}\subset B_{\epsilon_{*}}^{X}(x_{k})$ for $1\leq k\leq K$, and $\{h_{m}\}_{1\leq k\leq K}\subset E(X_{eq})$ such that
  • (1)
    $\tilde{\pi}(h_{k})\in\pi_{eq}(W_{k})$ for $k=1,2,\cdots,K$;
  • (2)
    $\tilde{\nu}(\cap_{k=1}^{K}\tilde{Z}_{*}h_{k}^{-1})>0.$
We can repeat the aforementioned process since the above items meets the conditions in Claim 4.5. By induction, there exist nonempty open sets $W_{a}\subset X$ for $a\in\bigcup_{n=1}^{\infty}\{1,\cdots,K\}^{n}$ and a sequence $(g_{n})_{n\in\mathbb{N}}$ of $G$ such that for every $n\in\mathbb{N}$ and $a=a_{1}a_{2}\cdots a_{n}\in\{1,\cdots,K\}^{n}$ the following conditions hold:
  • (1)
    $W_{ak}\subset W_{a}$ for $k\in\{1,\cdots,K\}$;
  • (2)
    $g_{n}W_{a}\subset B_{\epsilon_{*}}^{X}(x_{a_{n}})$;
  • (3)
    $g_{n}\notin\{g_{1},\cdots,g_{n-1}\}$.
Thus, for $N\in\mathbb{N}$ and $(a_{n})_{n=1}^{N}\in\{1,\cdots,K\}^{N}$, $\cap_{n=1}^{N}g_{n}^{-1}B_{\epsilon_{*}}^{X}(x_{a_{n}})\supset W_{a_{1}a_{2}\cdots a_{n}}\neq\emptyset$. This implies that $\{g_{n}\}_{n=1}^{\infty}$ is an infinite independence set for $(B_{\epsilon_{*}}^{X}(x_{1}),\cdots,B_{\epsilon_{*}}^{X}(x_{K}))$, and ends the proof of Theorem 4.1. ∎
At the end of this section, we prove Claim 4.5.
Proof of Claim 4.5.
Assume that nonempty open subsets $W_{1},\cdots,W_{M}$ of $X$, and $\{h_{m}\}_{1\leq m\leq M}\subset E(X_{eq})$ satisfy the following statements:
  • (P1)
    $\tilde{\pi}(h_{m})\in\pi_{eq}(W_{m})$ for $m=1,2,\cdots,M$;
  • (P2)
    $\tilde{\nu}(\cap_{m=1}^{M}(\tilde{Z}_{*}h_{m}^{-1}))>0.$
Given a finite subset $F$ of $G$, we aim to find $g_{*}\in G\setminus F$ and nonempty open sets
$W_{m,k}\subset W_{m}\text{ with }g_{*}W_{m,k}\subset B_{\epsilon_{*}}^{X}(x_{k})$
for $1\leq m\leq M,1\leq k\leq K$ and $\{h_{m,k}\}_{1\leq m\leq M,1\leq k\leq K}\subset E(X_{eq})$ such that
  • (P1’)
    $\tilde{\pi}(h_{m,k})\in\pi_{eq}(W_{m,k})$ for $m=1,2,\cdots,M$, $k=1,2,\cdots,K$;
  • (P2’)
    $\tilde{\nu}(\cap_{m=1}^{M}\cap_{k=1}^{K}(\tilde{Z}_{*}h_{m,k}^{-1}))>0.$
Indeed, by (P2) and Lemma 4.4, there exists an open neighborhood $\tilde{V}_{m}$ of $h_{m}$ for each $m=1,2,\cdots,M$ such that
$\tilde{\nu}(\cap_{m=1}^{M}\cap_{k=1}^{K}(\tilde{Z}_{*}h_{m,k}^{-1}))>0\text{ whenever }h_{m,k}\in\tilde{V}_{m}.$ (4.3)
Let
$W_{m}^{\prime}=\pi_{eq}^{-1}(\tilde{\pi}(\tilde{V}_{m}))\cap W_{m}.$ (4.4)
Since $\tilde{\pi}$ is an open map, it follows that $\tilde{\pi}(\tilde{V}_{m})$ is an open neighborhood of $\tilde{\pi}(h_{m})$. By (P1), $W_{m}^{\prime}\subset X$ is a nonempty open set. Since $\pi_{eq}$ is semi-open, we can find $r\in(0,\epsilon_{*})$ and $y_{m}\in X_{eq}$ such that
[…]
Denote
$\mathcal{N}=\{g\in G:B^{X_{eq}}_{r/2}(gy_{m})\cap Z_{*}\neq\emptyset,m=1,\cdots,M\}.$
For clarity, now we divide the remaining proof into three steps:
(S1): Prove $\mathcal{N}\in\mathcal{F}_{s}(G)$.
(S2): Prove there exist $g_{*}\in G\setminus F$, and nonempty open set $W_{m,k}\subset W_{m}^{\prime}$ such that $g_{*}W_{m,k}\subset B^{X}_{\epsilon_{*}}(x_{k})$ for $1\leq m\leq M,1\leq k\leq K$.
(S3): Prove there exists $h_{m,k}$ with $\tilde{\pi}(h_{m,k})\in\pi_{eq}(W_{m,k})$ for $1\leq m\leq M,1\leq k\leq K$ such that $\tilde{\nu}(\cap_{m=1}^{M}\cap_{k=1}^{K}(\tilde{Z}_{*}h_{m,k}^{-1}))>0$.
For (S1), denote $\tilde{V}_{m}^{\prime}=\tilde{\pi}^{-1}(B^{X_{eq}}_{r/2}(y_{m}))\cap\tilde{V}_{m}$. Then, $\tilde{V}_{m}^{\prime}\subset E(X_{eq})$ is an open set. By (4.5), $\tilde{\pi}(\tilde{V}_{m}^{\prime})=B^{X_{eq}}_{r/2}(y_{m})$, and hence $\tilde{V}_{m}^{\prime}$ is not empty. Using Theorem 4.3 (d), we know that $g\tilde{V}_{m}^{\prime}\cap\tilde{Z}_{*}\neq\emptyset$ if and only if $g\in\tilde{Z}_{*}\tilde{V}_{m}^{\prime-1}$. Thus,
$\begin{split}\mathcal{N}&\supset\{g\in G:g\tilde{V}_{m}^{\prime}\cap\tilde{Z}_{*}\neq\emptyset,m=1,\cdots,M\}\\ &=\{g\in G:g\in\tilde{Z}_{*}\tilde{V}_{m}^{\prime-1},m=1,\cdots,M\}\\ &=\{g\in G:g\in\cap_{m=1}^{M}\tilde{Z}_{*}\tilde{V}_{m}^{\prime-1}\}.\end{split}$ (4.6)
Since $E(X_{eq})$ is a topological group and $\tilde{V}_{m}^{\prime}\subset E(X_{eq})$ is an open set, it follows that $\tilde{V}_{m}^{\prime-1}\subset E(X_{eq})$ is an open set and so is $\tilde{Z}_{*}\tilde{V}_{m}^{\prime-1}$, for $m=1,2,\ldots,M$. Thus, $\cap_{m=1}^{M}\tilde{Z}_{*}\tilde{V}_{m}^{\prime-1}$ is an open subset of $E(X_{eq})$. Since $\tilde{V}_{m}^{\prime}\subset\tilde{V}_{m}$, by (4.3), one has
$\tilde{\nu}(\cap_{m=1}^{M}\tilde{Z}_{*}\tilde{V}_{m}^{\prime-1})>0.$
Now we deduce that $\cap_{m=1}^{M}\tilde{Z}_{*}\tilde{V}_{m}^{\prime-1}$ is a nonempty open subset of $E(X_{eq})$. Since $(E(X_{eq}),G)$ is a minimal tds (see Theorem 4.3 (b)), it follows from Lemma 2.2 (1) that
$\begin{split}\{g\in G:g\in\cap_{m=1}^{M}\tilde{Z}_{*}\tilde{V}_{m}^{\prime-1}\}\in\mathcal{F}_{s}(G).\end{split}$ (4.7)
Combining (4.6) and (4.7), it follows that $\mathcal{N}\in\mathcal{F}_{s}(G)$. Thus, we finish the proof of (S1).
For (S2), let
$\mathcal{F}_{m}=\{g\in G:\pi_{eq}^{-1}(B^{X_{eq}}_{r/2}(gy_{m}))\subset\overline{B^{X}_{r/2}(gW_{m}^{\prime})}\}\text{ for }m=1,\cdots,M.$
The equation (4.5) allows us to apply Lemma 3.9 (2) to derive
$\mathcal{F}_{m}\in\mathcal{F}_{ts}(G)\text{ for }m=1,\cdots,M.$
Since $\mathcal{N}\in\mathcal{F}_{s}(G)$, as established in (S1), it follows from Lemma 2.1 that
$\cap_{m=1}^{M}\mathcal{F}_{m}\cap\mathcal{N}\in\mathcal{F}_{s}(G).$
In particular, it is an infinite set, and so we can find $g_{*}\in(\cap_{m=1}^{M}\mathcal{F}_{m}\cap\mathcal{N})\setminus F.$ Since $g_{*}\in\mathcal{N}$, it follows that
$B^{X_{eq}}_{r/2}(gy_{m})\cap Z_{*}\neq\emptyset.$ (4.8)
Let $y_{m}^{\prime}\in B^{X_{eq}}_{r/2}(gy_{m})\cap Z_{*}$. Since $g_{*}\in\mathcal{F}_{m}$, it follows that $\pi_{eq}^{-1}(y_{m}^{\prime})\subset\overline{B^{X}_{r/2}(g_{*}W_{m}^{\prime})}.$ Thus, by the assumption $0<r<\epsilon_{*}$, we obtain that
[…]
Since $x_{1},x_{2},\cdots,x_{K}\in E_{*}$, it follows from (4.9) that
$B^{X}_{\epsilon_{*}}(x_{k})\cap g_{*}W_{m}^{\prime}\neq\emptyset,\text{ for $1\leq m\leq M,1\leq k\leq K$.}$
Thus, the set $W_{m,k}=(g_{*}^{-1}B^{X}_{\epsilon_{*}}(x_{k}))\cap W_{m}^{\prime}$ is a nonempty open subset of $X$ for $1\leq m\leq M,1\leq k\leq K$. It is clear that
$g_{*}W_{m,k}\subset B^{X}_{\epsilon_{*}}(x_{k})\text{ for $1\leq m\leq M,1\leq k\leq K$.}$
Therefore, (S2) is completed.
Finally, we prove (S3) to complete the proof of Claim 4.5. Denote $\tilde{W}_{m,k}=\tilde{\pi}^{-1}(\pi_{eq}(W_{m,k}))\cap\tilde{V}_{m}$. Let $h_{m,k}\in\tilde{W}_{m,k}$ for $1\leq m\leq M,1\leq k\leq K$. Hence,
$\tilde{\pi}(h_{m,k})\in\pi_{eq}(W_{m,k}),\text{ for $1\leq m\leq M,1\leq k\leq K$.}$
Since $\pi_{eq}(W_{m,k})\subset\pi_{eq}(W_{m}^{\prime})\subset\tilde{\pi}(\tilde{V}_{m}),$ one has $\tilde{W}_{m,k}\subset\tilde{V}_{m}$ is a nonempty set. Due to (4.3), we have
$\tilde{\nu}(\cap_{m=1}^{M}\cap_{k=1}^{K}(\tilde{Z}_{*}h_{m,k}^{-1}))>0.$
So, the proof is completed. ∎
5. Weakly mean-sensitive set
In this section, we prove Theorem 1.5 and present additional corollaries of Theorem 1.5.
5.1. Proof of Theorem 1.5 (i)
In this subsection, we establish Theorem 5.1, which together with Lemma 3.2, implies Theorem 1.5 (i). Meanwhile, we provide some corollaries of Theorem 5.1.
Theorem 5.1.
Let $(X,G)$ be a minimal tds with local Bronstein condition, where $G$ is amenable. Then each $E_{*}\in\mathcal{X}_{eq}^{\text{meas}}$ is a weakly mean-sensitive set.
Before proving the main theorem in this subsection, we establish the following lemma.
Lemma 5.2.
Let $G$ be amenable, and $F\in\mathcal{F}_{t}(G)$. Then there exists a Følner sequence $\{F_{n}\}_{n=1}^{\infty}$ of $G$ such that $\cup_{n\in\mathbb{N}}F_{n}\subset F$.
Proof.
By definition, it is straightforward to verify the fact that for any Følner sequence $\{\hat{F}_{n}\}_{\in\mathbb{N}}$ of $G$ and any sequence $\{g_{n}\}_{n\in\mathbb{N}}$ in $G$, $\{\hat{F}_{n}g_{n}\}_{\in\mathbb{N}}$ is also a Følner sequence.
Now we fix an Følner sequence $\{\hat{F}_{n}\}_{n=1}^{\infty}$ of $G$. Since $F\in\mathcal{F}_{t}(G)$, there exists $g_{n}\in G$ such that $\hat{F}_{n}g_{n}\subset F$ for each $n\in\mathbb{N}$. By the above fact, $\{\hat{F}_{n}g_{n}\}_{n=1}^{\infty}$ is the required Følner sequence. ∎
Now we prove Theorem 5.1.
Proof of Theorem 5.1.
Given $E_{*}\in\mathcal{X}_{eq}^{\text{meas}}$, let $K\in\mathbb{N}$ and $x_{1},x_{2},\cdots,x_{K}\in E_{*}$. Let $U_{k}$ be any open neighborhood of $x_{k}$, and $\epsilon_{*}>0$ be such that $B^{X}_{\epsilon_{*}}(x_{k})\subset U_{k}$ for each $k=1,2,\cdots,K$. To show $E_{*}$ is a weakly mean-sensitive set, it is sufficient to show that there exists $\delta>0$ such that for any nonempty open set $U\subset X$,
$\overline{BD}\{g\in G:(gU)\cap U_{k}\neq\emptyset\text{ for }k=1,2,\cdots,K\}\geq\delta.$
Since $E_{*}\in\mathcal{X}_{eq}^{\text{meas}}$, one has $Z_{*}:=\pi_{\mathcal{X}}(B_{\epsilon_{*}/2}^{\mathcal{X}}(E_{*}))$ has positive $\nu_{eq}$-measure and
$E_{*}\subset B^{X}_{\epsilon_{*}/2}(\pi^{-1}_{eq}(y))\text{ for every }y\in Z_{*}.$ (5.1)
We are going to show $\nu_{eq}(Z_{*})>0$ represents the desired $\delta$.
Now we fix a nonempty open set $U\subset X$. Since $\pi_{eq}$ is semi-open, we can find $y_{*}\in X_{eq}$ and $r\in(0,\frac{\epsilon_{*}}{2})$ such that $B_{r}^{X_{eq}}(y_{*})\subset\pi_{eq}(U)$. Since $(X_{eq},G)$ is an isometry, one has $B_{r/2}^{X_{eq}}(gy_{*})\cap Z_{*}\neq\emptyset$ if and only if $gy_{*}\in B^{X_{eq}}_{r/2}(Z_{*})$. Due to the unique ergodicity of $(X_{eq},G)$,
$\underline{BD}\{g\in G:B_{r/2}^{X_{eq}}(gy_{*})\cap Z_{*}\neq\emptyset\}=\underline{BD}\{g\in G:gy_{*}\in B^{X_{eq}}_{r/2}(Z_{*})\}\geq\nu_{eq}(Z_{*}).$ (5.2)
By Lemma 5.2 and Lemma 3.9 (2),
$\overline{BD}\{g\in G:\pi_{eq}^{-1}(B^{X_{eq}}_{r/2}(gy_{*}))\subset\overline{B^{X}_{r}(gU)}\}=1.$ (5.3)
Let $\mathcal{N}=\{g\in G:B_{r/2}^{X_{eq}}(gy_{*})\cap Z_{*}\neq\emptyset\text{ and}\pi_{eq}^{-1}(B^{X_{eq}}_{r/2}(gy_{*}))\subset\overline{B^{X}_{r}(gU)}\}$. Together with (5.2) and (5.3), one has
$\overline{BD}(\mathcal{N})\geq\nu_{eq}(Z_{*}).$ (5.4)
Notice that for every $g\in\mathcal{N}$, there exists $y(g)\in Z_{*}$ such that $\pi_{eq}^{-1}(y(g))\subset\overline{B^{X}_{r}(gU)}$. Since $r\in(0,\epsilon_{*}/2)$, the above implies
[…]
Noting that $x_{k}\in E_{*}$, one has for every $k\in\{1,2,\cdots,K\}$ and $g\in\mathcal{N}$,
$gU\cap U_{k}\supset gU\cap B_{\epsilon_{*}}(x_{k})\neq\emptyset.$
By (5.4), $(x_{1},\cdots,x_{K})$ is a weakly mean-sensitive tuple, and so $E_{*}$ is a weakly mean-sensitive set. ∎
The following corollary can be proved using exactly the same argument as in Corollary 4.2, with Theorem 4.1 replaced by Theorem 5.1. So we omit the proof.
Corollary 5.3.
Let $(X,G)$ be a minimal tds with local Bronstein condition, where $G$ is amenable. Then, the following statements hold:
  1. (i)
    If $(X,G)$ has no essential weakly mean-sensitive $K$-tuple for some $K\geq 2$, then $(X,G)$ is regular $K^{\prime}$ to one for some $11\leq K^{\prime}\leq K-$ 1.
  2. (ii)
    If $\pi_{eq}$ is open, then $\pi_{eq}^{-1}(y)$ is a weakly mean-sensitive set for every $y\in X_{eq}$.
  3. (iii)
    Let $N\in\mathbb{N}\cup\{\infty\}$ and $K\in\mathbb{N}$ with $2\leq K\leq N$. If $\pi_{eq}$ is almost $N$ to one, then for every $y\in X_{eq}$, $\pi_{eq}^{-1}(y)$ contains essential weakly mean-sensitive $K$-tuples.
  4. (iv)
    If $(X,G)$ has no essential weakly mean-sensitive $K$-tuple for some $K\geq 2$, then $|\mathcal{M}^{e}(X,G)|\leq K-1$.
5.2. Proof of Theorem 1.5 (ii) and (iii)
To prove Theorem 1.5 (ii) and (iii), we need a lemma.
Lemma 5.4.
Let $(X,G)$ be a minimal tds, where $G$ is amenable. If a $K$-tuple $(x_{1},\dots,x_{K})\in X^{K}$ is a weakly mean-sensitive tuple, then for any open neighborhood $U_{k}$ of $x_{k}$ for $k=1,2,\cdots,K$, one has $\nu_{eq}(\cap_{k=1}^{K}\pi_{eq}(U_{k}))>0$.
Proof.
Choose an open neighborhood $U_{k}^{\prime}$ of $x_{k}$ with $\overline{U_{k}^{\prime}}\subset U_{k}$ for $k=1,2,\cdots,K$. By definition of weakly mean-sensitive tuple, there is $\delta_{*}>0$ such that for any nonempty open set $U\subset X$,
$\overline{BD}\{g\in G:U_{k}^{\prime}\cap gU\neq\emptyset\text{ for }k=1,2\cdots,K\}\geq\delta_{*}.$
This implies that for any nonempty open set $V\subset X_{eq}$,
$\overline{BD}\{g\in G:\pi_{eq}(U_{k}^{\prime})\cap gV\neq\emptyset\text{ for }k=1,2\cdots,K\}\geq\delta_{*}.$ (5.5)
We will prove that $\nu_{eq}(\cap_{k=1}^{K}\pi_{eq}(U_{k}))\geq\delta_{*}>0$. Suppose, for the sake of contradiction, that $\nu_{eq}(\cap_{k=1}^{K}\pi_{eq}(U_{k}))<\delta_{*},$ and hence
$\nu_{eq}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}}))<\delta_{*}.$ (5.6)
Noting that $\pi_{eq}(x_{1})=\cdots=\pi_{eq}(x_{K})$, $\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}})$ is not empty. Since $\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}})$ is compact and nonempty, by (5.6), and a standard argument, there exists $\epsilon>0$ such that
$\nu_{eq}\left(\overline{B_{\epsilon}^{X_{eq}}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}}))}\right)<\delta_{*}.$ (5.7)
Claim 5.5.
There exists $\epsilon_{*}>0$ such that for every $\epsilon_{*}^{\prime}\in(0,\epsilon_{*})$ and $y\in X_{eq}$, if $B_{\epsilon_{*}^{\prime}}^{X_{eq}}(y)\cap\pi_{eq}(\overline{U_{k}^{\prime}})\neq\emptyset$ for $k=1,2,\cdots,K$, then
$B_{\epsilon_{*}^{\prime}}^{X_{eq}}(y)\cap B_{\epsilon}^{X_{eq}}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}}))\neq\emptyset.$
Proof of Claim 5.5.
If Claim 5.5 does not hold, we can find $\epsilon_{n}\searrow 0$ as $n\to\infty$ and $\{y_{n}\}_{n\in\mathbb{N}}\subset X_{eq}$ such that $B_{\epsilon_{n}}^{X_{eq}}(y_{n})\cap\pi_{eq}(\overline{U_{k}^{\prime}})\neq\emptyset$ for $k=1,2,\cdots,K$ and
$B_{\epsilon_{n}}^{X_{eq}}(y_{n})\cap B_{\epsilon}^{X_{eq}}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}}))=\emptyset.$ (5.8)
Note $B_{k}=\pi_{eq}(\overline{U_{k}^{\prime}})\setminus B_{\epsilon}^{X_{eq}}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}}))$ for $k=1,2,\cdots,K$. By (5.8), $B_{\epsilon_{n}}^{X_{eq}}(y_{n})\cap\pi_{eq}(\overline{U_{k}^{\prime}})\subset B_{k}$. Choose $y_{n}^{(k)}\in B_{\epsilon_{n}}^{X_{eq}}(y_{n})\cap\pi_{eq}(\overline{U_{k}^{\prime}})\subset B_{k}$ for $n\in\mathbb{N}$ and $k=1,2,\cdots,K$. Passing by a subsequence, we assume that $y_{n}^{(k)}\to y^{(k)}$ as $n\to\infty$ for $k=1,2,\cdots,K$. Since $y_{n}^{(k)}\in B_{k}$ and $B_{k}$ is compact, one has $y^{(k)}\in B_{k}$ for $k=1,2,\cdots,K$. Noting that $\epsilon_{n}\searrow 0$ as $n\to\infty$ and $\{y_{n}^{(k)}:k=1,2,\cdots,K\}\subset B_{\epsilon_{n}}^{X_{eq}}(y_{n})$, one has
$y^{(1)}=y^{(2)}=\cdots=y^{(K)}.$
Thus, $\cap_{k=1}^{K}B_{k}\neq\emptyset$, contradicting its definition. Hence, Claim 5.5 holds. ∎
Let $\epsilon_{*}>0$ be as in Claim 5.5. By (5.7) and a standard argument, we can find $\epsilon_{*}^{\prime}\in(0,\epsilon_{*})$ such that
$\nu_{eq}(\partial(B_{\epsilon_{*}^{\prime}+\epsilon}^{X_{eq}}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}}))))=0,$ (5.9)
and
$\nu_{eq}(B_{\epsilon_{*}^{\prime}+\epsilon}^{X_{eq}}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}})))<\delta_{*}.$ (5.10)
Fixing $y_{*}\in X_{eq}$, by the unique ergodicity of $(X_{eq},G)$, (5.9) and (5.10), we have
$\overline{BD}\{g\in G:gy_{*}\in B_{\epsilon_{*}^{\prime}+\epsilon}^{X_{eq}}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}}))\}<\delta_{*}.$ (5.11)
By Claim 5.5, if $B^{X_{eq}}_{\epsilon_{*}^{\prime}}(y)\cap\pi_{eq}(\overline{U_{k}^{\prime}})\neq\emptyset$ for all $k=1,2,\cdots,K$, then
$B_{\epsilon_{*}^{\prime}}^{X_{eq}}(y)\cap B_{\epsilon}^{X_{eq}}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}}))\neq\emptyset,$
and thus
$y\in B^{X_{eq}}_{\epsilon_{*}^{\prime}}(y)\subset B_{\epsilon_{*}^{\prime}+\epsilon}^{X_{eq}}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}})).$
Combining this with (5.11) and the fact $(X_{eq},G)$ is an isometry, one has
$\overline{BD}\{g\in G:gB^{X_{eq}}_{\epsilon_{*}^{\prime}}(y_{*})\cap\pi_{eq}(\overline{U_{k}^{\prime}})\neq\emptyset\text{ for }k=1,2,\cdots,K\}$
$=\overline{BD}\{g\in G:B^{X_{eq}}_{\epsilon_{*}^{\prime}}(gy_{*})\cap\pi_{eq}(\overline{U_{k}^{\prime}})\neq\emptyset\text{ for }k=1,2,\cdots,K\}$
$\leq\overline{BD}\{g\in G:gy_{*}\in B_{\epsilon_{*}^{\prime}+\epsilon}^{X_{eq}}(\cap_{k=1}^{K}\pi_{eq}(\overline{U_{k}^{\prime}}))\}$
$<\delta_{*}.$
This is in contradiction to (5.5) (by taking $V=B^{X_{eq}}_{\epsilon_{*}^{\prime}}(y_{*})$), completing the proof of Lemma 5.4. ∎
Now we prove Theorem 1.5 (ii).
Proof of Theorem 1.5 (ii).
Assume that $\pi_{eq}$ is regular $K$ to one. By Theorem 1.5 (i), $(X,G)$ has essential weakly mean-sensitive $K$-tuples. Now we show that $(X,G)$ does not have essential weakly mean-sensitive $(K+1)$-tuples.
By contrary, we suppose that $(x_{1},\cdots,x_{K+1})$ is an essential weakly mean-sensitive $(K+1)$-tuple. We choose $\epsilon>0$ sufficiently small such that $B^{X}_{\epsilon}(x_{1}),\cdots,B^{X}_{\epsilon}(x_{K+1})$ are pairwise disjoint. By Lemma 5.4,
$\nu_{eq}(\cap_{k=1}^{K+1}\pi_{eq}(B^{X}_{\epsilon}(x_{k})))>0.$ (5.12)
Since $\pi_{eq}$ is regular $K$ to one, one has
$\nu_{eq}(\{y\in X_{eq}:|\pi_{eq}^{-1}(y)|=K\})=1.$
Together this with (5.12), we can choose $y_{*}\in\cap_{k=1}^{K+1}\pi_{eq}(B^{X}_{\epsilon}(x_{k}))$ such that $|\pi_{eq}^{-1}(y_{*})|=K$ and $B^{X}_{\epsilon}(x_{k})\cap\pi_{eq}^{-1}(y_{*})\neq\emptyset$ for $k\in\{1,2,\cdots,K+1\}.$ By pigeonhole principle there exists $11\leq k_{1}<k_{2}\leq K+$ 1 such that
$B^{X}_{\epsilon}(x_{k_{1}})\cap B^{X}_{\epsilon}(x_{k_{2}})\cap\pi_{eq}^{-1}(y_{*})\neq\emptyset,$
which contradicts the assumption $B^{X}_{\epsilon}(x_{1}),\cdots,B^{X}_{\epsilon}(x_{K+1})$ are pairwise disjoint. Now the proof is completed. ∎
Finally, we prove Theorem 1.5 (iii).
Proof of Theorem 1.5 (iii).
Let $E_{*}$ be a weakly mean-sensitive set of $(X,G)$. Let $K\in\mathbb{N}$ and $x_{1},\cdots,x_{K}\in E_{*}$. Let $U_{k}$ be an open neighborhood of $x_{k}$ for $k=1,2,\cdots,K$. By Lemma 5.4,
$\nu_{eq}(\cap_{k=1}^{K}\pi_{eq}(U_{k}))>0.$
By Theorem 1.2, there exists $\tilde{y}\in\cap_{k=1}^{K}\pi_{eq}(U_{k})$ such that $\pi^{-1}_{eq}(\tilde{y})$ is an IT-set. By taking $\tilde{x}_{k}\in U_{k}\cap\pi^{-1}_{eq}(\tilde{y})$ for $k=1,2,\cdots,K$, we have a $K$-IT-tuple $(\tilde{x}_{1},\cdots,\tilde{x}_{K})$, and so $(U_{1},\cdots,U_{K})$ has infinite independence sets. Thus, by the arbitrariness of $(U_{1},\cdots,U_{K})$, one has $(x_{1},\cdots,x_{K})$ is an IT-tuple, and so $E_{*}$ is an IT-set. ∎
Acknowledgment. L. Xu is partially supported by National Key R&D Program of China (No. 2024YFA1013602, 2024YFA1013600) and National Natural Science Foundation of China grants (12031019, 12371197, 12426201).